### v1

# Rethinking Layer Removal: Preserving Critical Components with Task-Aware Singular Value Decomposition

Abstract
Layer removal has emerged as a promising approach for compressing large language models (LLMs) by leveraging redundancy within layers to reduce model size and accelerate inference. However, this technique often compromises internal consistency, leading to performance degradation and instability, with varying impacts across different model architectures. In this work, we propose Taco-SVD, a task-aware framework that retains task-critical singular value directions, preserving internal consistency while enabling efficient compression. Unlike direct layer removal, Taco-SVD preserves task-critical transformations to mitigate performance degradation. By leveraging gradient-based attribution methods, Taco-SVD aligns singular values with downstream task objectives. Extensive evaluations demonstrate that Taco-SVD outperforms existing methods in perplexity and task performance across different architectures while ensuring minimal computational overhead.
[…]
Rethinking Layer Removal: Preserving Critical Components with Task-Aware Singular Value Decomposition
Kainan Liu1,2,†, Yong Zhang1,†, Ning Cheng1,*, Zhitao Li1, Shaojun Wang1, Jing Xiao1, 1Ping An Technology (Shenzhen) Co., Ltd., China 2The Hong Kong University of Science and Technology (Guangzhou) {zhangyong203, chengning211}@pingan.com.cn
footnotetext: † Equal contribution.footnotetext: *Corresponding author.footnotetext: This work was done during Kainan Liu’s internship at Ping An Technology.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing tasks, including language generation, reasoning, and question answering (Brown et al., 2020; Touvron et al., 2023). However, their massive parameter sizes create significant computational and memory challenges, making them difficult to deploy on resource-constrained devices (Zhou et al., 2024). To address this, model compression techniques such as quantization (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023), knowledge distillation (Gu et al., 2023; Xu et al., 2024), and pruning(Sun et al., 2024; Ashkboos et al., 2024) have been widely explored. Among these, structured pruning methods, especially layer removal (Men et al., 2024; Yang et al., 2024; Song et al., 2024), have shown promise by leveraging redundancy within layers of LLMs to achieve significant parameter reductions and inference speedup.
Figure 1: Zero-shot performance of Taco-SVD and pure layer removal under 20% compression ratio. Dashed lines represent perplexity on WikiText-2, while bar charts show the percentage of the original model’s average accuracy on seven reasoning datasets.
Despite its simplicity and computational efficiency, pure layer removal introduces two notable challenges, as shown in Figure 1. First, while task accuracy may remain stable in certain cases, perplexity (PPL) often degrades significantly, reflecting disruptions to the model’s internal consistency, which stems from impaired information flow and misaligned intermediate representations (Liang et al., 2024). This increase in PPL suggests that layer removal may affect critical transformations responsible for maintaining coherent representations, potentially reducing the model’s generalization ability on complex or unseen tasks.
Second, the impact of layer removal varies significantly across model architectures. For instance, under the same pruning ratio, LLaMA2 (Touvron et al., 2023) maintains reasonable performance, whereas LLaMA3 (Dubey et al., 2024) experiences a steep performance drop. This observation indicates that the tolerance to layer pruning is architecture-dependent. Such variability underscores a lack of robustness in pure layer removal, where certain task-critical transformations may be unevenly affected, resulting in performance inconsistencies across different architectures.
Our analysis suggests that these issues arise from the removal of transformations that play an essential role in maintaining both internal consistency and task-specific performance. While existing methods, such as LoRA-based fine-tuning (Hu et al., 2022) or lightweight layer replacements (Chen et al., 2024), aim to recover or approximate the lost functionality, they often incur substantial computational costs and do not fully address the challenge of preserving task-critical transformations in a resource-efficient manner.
To address these challenges, we propose Taco-SVD (Task-Aware Singular Value Decomposition for LLM Compression), a novel framework designed to address the limitations of pure layer removal by preserving task-critical transformations. Taco-SVD fundamentally rethinks layer removal by integrating task-aware optimization into singular value decomposition, ensuring that compression decisions align with task-specific objectives. Unlike conventional SVD, which focuses on global matrix reconstruction, Taco-SVD leverages gradient attribution to directly quantify the task-specific importance of singular values. Using a small calibration dataset from a general-purpose, unlabeled source, our method identifies and preserves the most impactful transformations within redundant layers. This approach minimizes the indiscriminate loss of critical task-specific transformations, maintaining internal consistency and representation flow while achieving efficient compression.
To validate our approach, we conducted experiments on tasks such as language generation, commonsense reasoning, and mathematical reasoning using LLaMA and Mistral models. Taco-SVD achieves parameter reduction comparable to pure layer removal while improving perplexity and task performance. Notably, Taco-SVD demonstrates consistent robustness across diverse model architectures, maintaining strong performance without additional training. It also achieves strong performance with minimal calibration data—only 64 samples from a general-purpose, unlabeled dataset, demonstrating its effectiveness. Moreover, fine-tuning further enhances its performance, highlighting Taco-SVD’s robustness and efficiency in compressing large language models.
Figure 2: Overview of Taco-SVD. The figure illustrates how Taco-SVD uses gradient attribution to identify task-critical singular values and applies SVD to compress redundant layers while preserving task-relevant transformations.
2 Related Work
2.1 Layer Removal Techniques and Their Impact on Model Consistency
Layer removal is a structured pruning technique that eliminates redundant layers within Large Language Models (LLMs), aiming to reduce the model’s computational and memory requirements. Methods such as ShortGPT (Men et al., 2024) introduce a metric called Block Influence to assess the significance of individual layers, enabling efficient one-shot removal of less important layers. SLEB (Song et al., 2024) improves this by employing an iterative pruning strategy, evaluating the importance of each layer based on the current state of the layer-removed LLMs. LaCo (Yang et al., 2024), on the other hand, proposes a gradual compression approach, progressively merging redundant layers from deeper to shallower parts of the network.
Although effective in reducing model size, layer removal often disrupts internal consistency, leading to degraded task performance and increased perplexity, as shown in our analysis building on the concept introduced by Liang et al. (2024). Additionally, its sensitivity to different architectures limits generalizability, with some models tolerating pruning better than others. These challenges highlight the need for approaches that preserve task-critical information while ensuring robust compression across diverse scenarios.
2.2 Post-Pruning Compensation and Training-Based Recovery Approaches
To mitigate the degradation caused by layer removal, post-pruning compensation methods have been developed. (Kim et al., 2024) introduced Shortened LLaMA, which incorporates LoRA (Hu et al., 2022) to recover the ability of layer removal models. However, LoRA alters the original singular value spectrum, which may diminish the contribution of pre-trained features (Shuttleworth et al., 2024). To address this issue, LLM-Streamline (Chen et al., 2024) trains a lightweight module such as feed-forward neural network (FFN) or a transformer layer to approximate the functionality of the pruned redundant layers. Although effective to some extent, these methods incur significant computational and data costs, limiting their practicality in resource-constrained scenarios.
2.3 Singular Value Decomposition for LLM Compression
Singular Value Decomposition (SVD) is a foundational technique for compressing dense models by decomposing weight matrices into low-rank structures. However, vanilla SVD focuses solely on global matrix reconstruction, providing general approximations without considering task-specific characteristics. Recent methods have enhanced SVD to better align with model and task requirements. FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD. ASVD (Yuan et al., 2023) uses activation patterns from a calibration dataset to scale weight matrices, reducing compression-induced activation errors. SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss. Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices.
Despite these advancements, existing methods mainly focus on optimizing compression loss without directly addressing the relationship between singular values and task-specific losses. Taco-SVD bridges this gap by integrating task-aware optimization into SVD, leveraging gradient attribution to retain task-critical singular values and align compressed models with task objectives.
3 Method
3.1 Overview and Motivation
In this paper, we propose a novel task-aware compression framework named Taco-SVD, designed to compress redundant layers in large language models (LLMs) while preserving task-critical transformations. Unlike traditional layer removal methods that directly eliminate redundant components, Taco-SVD integrates task-aware singular value decomposition into the layer removal paradigm. By introducing gradient-based attributions, it systematically aligns the importance of singular values with task-specific objective, ensuring that the most essential components within redundant layers are retained during truncation.
Figure 2 provides an overview of Taco-SVD. The method consists of two main steps: (1) identifying redundant layers based on cosine similarity, and (2) applying task-aware singular value decomposition (SVD) to retain the most critical singular values and their corresponding singular vectors. Below, we describe each step in detail.
3.2 Redundant Layer Selection
The first step in Taco-SVD is identifying redundant layers. These are layers that contribute minimally to the transformation of hidden states, exhibiting high redundancy and limited impact on overall model performance. Following prior works (Song et al., 2024; Chen et al., 2024), we use cosine similarity to quantify the degree of transformation in a given layer.
For a transformer layer with input hidden state $H_{i}\in\mathbb{R}^{d}$ and output hidden state $H_{i+1}\in\mathbb{R}^{d}$, the cosine similarity is computed as:
$\text{cos}(H_{i},H_{i+1})=\frac{H_{i}^{T}H_{i+1}}{\|H_{i}\|_{2}\|H_{i+1}\|_{2}}$ (1)
A high cosine similarity indicates minimal transformation, suggesting that the layer is redundant. Instead of directly removing these layers, Taco-SVD compresses their weight matrices using a task-aware approach to preserve task-relevant transformations.
3.3 Task-Aware Singular Value Decomposition
3.3.1 Motivation for Task-Aware SVD
Singular value decomposition (SVD) is a standard technique for compressing weight matrices by providing a global low-rank approximation. Traditional SVD methods retain the top-$k$ singular values based solely on their magnitudes, which are task-agnostic and often fail to capture task-relevant components. This can lead to the removal of critical transformations necessary for downstream tasks (Hsu et al., 2022; Wang et al., 2024).
To address this limitation, Taco-SVD introduces task awareness into SVD by using gradient-based attribution to evaluate the task-specific importance of each singular value and its corresponding singular vectors. This ensures that only the most task-relevant transformations are preserved.
3.3.2 Singular Value Decomposition (SVD)
For a redundant layer’s weight matrix $W\in\mathbb{R}^{m\times n}$, SVD decomposes $W$ as:
$W=U\Sigma V^{T}$ (2)
where $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{k})$ contains the singular values ($\sigma_{i}\geq 0$), and $U\in\mathbb{R}^{m\times m}$, $V\in\mathbb{R}^{n\times n}$ are orthogonal matrices containing the left and right singular vectors, respectively.
3.3.3 Task-Aware Singular Value Importance
To assess the task-specific importance of each singular value, we introduce $\Phi(\sigma_{i})$, the Task-Aware Singular Value Importance, which quantifies the contribution of $\sigma_{i}$ to the task loss $\mathcal{L}_{\text{task}}$. Using its first-order Taylor expansion, the alteration in the task loss $L$ caused by $\sigma_{i}$ is approximated as:
$\Delta\mathcal{L}(\sigma_{i})\approx\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}\Delta\sigma_{i}$ (3)
where $\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}$ represents the gradient sensitivity of $\sigma_{i}$ to the task loss, and $\Delta\sigma_{i}$ is the perturbation applied to $\sigma_{i}$.
The gradient sensitivity is calculated by projecting the gradient of $\mathcal{L}_{\text{task}}$ with respect to $W$ onto the singular directions:
$\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}=\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$ (4)
where $G=\frac{\partial\mathcal{L}_{\text{task}}}{\partial W}$ is the gradient of the task loss with respect to $W$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the left and right singular vectors corresponding to $\sigma_{i}$. The detailed proof can be found in the Appendix A.1.
Integrating the gradient sensitivity with the magnitude of $\sigma_{i}$, we define $\Phi(\sigma_{i})$ as:
$\Phi(\sigma_{i})=\sigma_{i}\cdot\left(\mathbf{u}_{i}^{T}G\mathbf{v}_{i}\right)$ (5)
where $\Phi(\sigma_{i})$ measures the overall contribution of $\sigma_{i}$ to task performance, considering both its magnitude and alignment with the task gradient. Larger value of $\Phi(\sigma_{i})$ indicates greater importance.
We rank singular values in descending order of task relevance based on $\Phi(\sigma_{i})$. Only the top-$k$ singular values with the highest $\Phi(\sigma_{i})$ are retained, and the compressed weight matrix is reconstructed as:
$W_{\text{task}}=\sum_{i=1}^{k}\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{T}$ (6)
This ensures that $W_{\text{task}}$ preserves the transformations most critical to task performance, achieving efficient compression while maintaining task consistency.
3.4 Layer-wise Low-Rank Decomposition
Taco-SVD applies a layer-wise backward compression strategy, starting from the last redundant layer and moving sequentially backward. For each redundant layer, weight matrices are compressed using Task-Aware Singular Value Decomposition, retaining only the top-$k$ singular values ranked by $\Phi(\sigma_{i})$. This ensures a low-rank approximation that preserves task-relevant transformations.
By processing one layer at a time, this approach significantly reduces memory overhead, avoiding the need to load the entire model into memory. Additionally, the backward order ensures that compressing deeper layers does not disrupt the consistency of earlier layers, preserving the overall stability of the model. The detailed implementation for Layer-wise Low-Rank Decomposition is provided in Algorithm 1 in the Appendix A.2.
4 Experiments
4.1 Experimental Setup
To validate the effectiveness of Taco-SVD, we conducted experiments across multiple large language models (LLMs) and datasets, focusing on evaluating task performance, cross-model robustness, and computational efficiency. Below, we detail the models, datasets, baselines, and experimental configurations.
Models
We evaluated our method on a diverse set of LLMs including LLaMA family models LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023), LLaMA3.1-8B-Instruct (Dubey et al., 2024) and Mistral-7b (Jiang et al., 2023). These models enable a comprehensive assessment of our method across varying sizes and architectures, highlighting its generability and robustness.
Datasets
Ten datasets are used as evaluation benchmarks including three language generation datasets: WikiText-2 (Merity et al., 2017), PTB (Marcus et al., 1993), C4 (Raffel et al., 2020) and seven datasets for common sense and mathematical reasoning include: WinoGrande (Sakaguchi et al., 2020), Hellaswag (Zellers et al., 2019), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-c, ARC-e (Clark et al., 2018) for common sense reasoning, and MathQA (Amini et al., 2019) for mathematical reasoning.
Metric
We use perplexity and accuracy for language generation tasks and reasoning tasks, respectively. All reasoning tasks are tested in zero-shot setting using the LM-Evaluation-Harness framework (Gao et al., 2024).
Baselines
We compare Taco-SVD against training-free and training-based compression methods at the same compression ratio. For training-free methods, we select the layer removal technique LaCo (Yang et al., 2024), as well as a structured pruning method SliceGPT (Ashkboos et al., 2024) and the SVD-based LLM compression method ASVD (Yuan et al., 2023). For training-based methods, we compare Taco-SVD with LoRA (Hu et al., 2022) and the layer replacement method LLM-Streamline (Chen et al., 2024) with two variants: LLM-Streamline-FFN and LLM-Streamline-Layer, which replace redundant layers with Feed-Forward Neural Network (FFN) and Transformer Layer (Layer), respectively.
Implementation Details
To ensure a fair comparison, all experimental setups are consistent across all methods. Specifically, for training-free methods, we randomly select 512 samples from the WikiText-2 dataset as calibration data. For training-based methods, all models are fine-tuned on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32. We use AdamW (Loshchilov and Hutter, 2019) as our optimizer and set the learning rate to $3\times 10^{-4}$. All our experiments are conducted on a single NVIDIA A100 GPU with mixed precision enabled. Further experimental details can be found in the Appendix A.3.
Model Method Ratio WikiText-2 PTB C4 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Mistral-7B Dense 0% 5.25 23.73 8.36 32.6 80.85 73.88 61.22 50.43 80.58 35.88 59.35
LaCo 21.1% 244.53 524.26 243.98 20.2 35.17 57.54 26.23 24.66 52.77 24.29 34.41
SliceGPT 20% 9.04 107.65 38.17 19.2 51.05 59.43 34.84 25.26 60.72 23.25 39.11
ASVD 20% 19.28 171.54 27.93 19.8 55.68 55.96 35.84 24.49 66.21 24.19 40.31
Taco-SVD 20.3% 26.42 77.98 28.88 21.4 51.47 68.19 42.82 35.24 66.59 25.83 44.51
LLaMA2-7B Dense 0% 5.47 20.83 7.29 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
LaCo 18.1% 39.43 93.98 71.77 26.2 48.36 59.19 41.93 32.42 68.72 24.02 42.98
SliceGPT 21.5% 8.63 99.25 40.7 22 54.34 60.62 37.4 27.99 63.38 23.18 41.27
ASVD 20% 9.7 88.39 12.46 27.8 58.41 61.09 46.97 32.42 71.05 25.09 46.12
Taco-SVD 21.6% 16.12 44.09 23.13 24 53.59 62.9 42.94 32.76 71.06 22.88 44.30
LLaMA-2-13B Dense 0% 4.88 28.92 6.99 32.4 73.27 69.61 59.72 45.56 78.78 30.02 55.62
LaCo 19.5% 27.26 59.14 27.42 27.6 52.25 62.98 42.52 33.36 69.91 25.32 44.85
SliceGPT 21.5% 7.44 99.77 37.65 29.4 58.92 64.8 38.96 31.74 64.42 23.65 44.56
ASVD 20% 8.07 101.1 11.86 28.8 63.55 62.59 46.12 32.68 73.72 24.65 47.44
Taco-SVD 20.6% 11.1 56.57 14.76 26.4 60.62 66.22 46.55 35.07 73.11 24.22 47.46
LLaMA3.1-8B -Instruct Dense 0% 7.21 12.33 12.06 33.8 81.9 73.72 59.1 51.71 79.87 39.46 59.94
LaCo 19% 753.98 1577.44 807.12 26 48.78 64.56 33.2 29.61 65.07 30.25 42.50
SliceGPT 20% 52.78 460.73 324.71 14.6 43.48 51.38 29.6 22.78 58.38 21.78 34.57
ASVD 20% 2443.99 4469.54 2842.23 11.8 28.37 48.62 26.63 19.2 54.84 18.79 29.75
Taco-SVD 20.1% 37.86 63.97 42.25 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
Table 1: Zero-shot performance of Taco-SVD and training-free baselines under 20% compression ratio. Results are reported on three language modeling datasets (perplexity) and seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
Model Method Ratio WikiText-2 PTB C4 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Mistral-7B Dense 0% 5.25 23.73 8.36 32.6 80.85 73.88 61.22 50.43 80.58 35.88 59.35
LoRA 25% 11.56 34.05 14.86 24.6 68.73 70.72 51.18 40.1 71.93 26.37 50.52
LS-FFN 25.2% 16.12 59.98 18.4 21.8 66.88 68.03 49.92 37.8 70.35 27.2 48.85
LS-Layer 24.1% 14.67 54.31 16.61 21.6 68.77 70.09 45.02 39.25 72.25 26.83 49.12
Taco-SVD 24.3% 11.62 44.14 14.85 26.2 69.28 70.32 51.7 40.87 73.07 27.17 51.23
LLaMA-2-7B Dense 0% 5.47 20.83 7.29 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
LoRA 25% 11.7 35.39 13.95 26.2 57.79 63.06 48.15 33.28 71.16 24.15 46.26
LS-FFN 25.7% 13.94 37.43 14.61 24.6 55.98 62.51 48.32 33.28 71 25.19 45.84
LS-Layer 24% 11.57 32.62 12.91 25.4 59.85 61.88 50.21 35.15 72.91 23.85 47.04
Taco-SVD 24.2% 9.59 30.38 12.22 28.4 60.52 64.56 50.32 36.01 71.76 24.85 48.06
LLaMA-2-13B Dense 0% 4.88 28.92 6.99 32.4 73.27 69.61 59.72 45.56 78.78 30.02 55.62
LoRA 25% 8.46 58.15 11.35 26.6 63.05 67.4 53.47 38.4 73.5 25.33 49.68
LS-FFN 25.4% 8.11 76.3 11.22 28.4 61.07 67.17 54.04 38.23 73.29 25.49 49.67
LS-Layer 24.4% 8.23 77.04 11.09 29 62.42 67.25 54.25 39.76 74.21 26.06 50.42
Taco-SVD 24.2% 7.52 38.31 10.42 28.6 63.3 67.96 54.63 38.74 73.88 27.27 50.63
LLaMA3.1-8B -Instruct Dense 0% 7.21 12.33 12.06 33.8 81.9 73.72 59.1 51.71 79.87 39.46 59.94
LoRA 22.3% 14.56 21.24 21.8 27.6 69.65 69.38 50.54 41.3 72.69 30.18 51.62
LS-FFN 22% 16.79 24.76 24.14 26.4 70.24 70.17 48.97 40.53 72.03 30.25 51.23
LS-Layer 21.7% 14.27 20.85 20.79 27.2 73.53 69.38 51.22 42.66 73.39 31.56 52.71
Taco-SVD 21% 14.13 20.99 21.73 28.2 72.01 69.93 51.5 43.26 73.5 31.26 52.81
Table 2: Performance comparison of Taco-SVD + FT and training-based baselines under 25% compression ratio. LS-FFN and LS-Layer denote two variants of LLM-Streamline. Results are reported on three language modeling datasets (perplexity) and seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
4.2 Results
We conducted comprehensive experiments to evaluate Taco-SVD from four key perspectives: (a) zero-shot compression performance on language generation and reasoning tasks, (b) effectiveness as an initialization method compared to training-based approaches, (c) robustness at high compression ratios, and (d) compression efficiency and inference speed on real hardware.
Training-Free Compression
To evaluate the generalizability of Taco-SVD, we compared its performance with training-free baselines under approximately 20% compression ratio across four models with varying scales and architectures: Mistral-7B, LLaMA2-7B, LLaMA2-13B, and LLaMA3.1-8B-Instruct. The results, summarized in Table 1, demonstrate that Taco-SVD consistently achieves strong performance across all models and tasks, maintaining competitive perplexity and accuracy even on challenging architectures like LLaMA3. Notably, Taco-SVD exhibits significantly improved robustness compared to baselines, effectively addressing the variability in sensitivity to pruning across diverse architectures.
Training-based Compression
In this section, we further demonstrate that Taco-SVD can provide better initialization for fine-tuning, enabling faster and more accurate recovery of the compressed model compared to other methods. To ensure that the number of trainable parameters remains approximately the same across all methods, we retain only 10% of the parameters in each redundant layer and set only the parameters within these redundant layers to be trainable. As shown in Table 2, Taco-SVD consistently outperforms LoRA and LLM-Streamline across different LLMs. In addition, we plot the validation loss curves during the training process for each method in the Appendix A.4, we can observe that Taco-SVD has better initialization, leading to more stable and faster convergence during fine-tuning.
Performance at High Compression Ratio
To evaluate the performance and stability of our proposed method at high compression ratios, we compare Taco-SVD with the baselines on LLaMA2-7B at compression ratios of 25%, 30%, 35% and 40%. The perplexity (PPL) on WikiText-2, PTB and the average accuracy across all reasoning benchmarks are shown in Figure 3, with detailed results for each benchmark presented in the Appendix A.5. We make the following three observations: (a) Most layer removal methods cannot tolerate high compression ratios, with performance dropping sharply at higher compression levels. (b) Compared to training-free baselines, Taco-SVD consistently achieves better accuracy across all compression ratios. (c) Fine-tuning Taco-SVD significantly recovers the accuracy of compressed models, even at high compression ratios.
Figure 3: Zero-shot performance of LLaMA2-7B compressed by Taco-SVD and baselines under 20% to 40% compression ratio. Top: Perplexity on WikiText-2 (left) and PTB (right). Bottom: Average accuracy on seven reasoning benchmarks.
Compression Costs and Acceleration Effect
Taco-SVD can not only compress LLMs at very low cost but also improve inference efficiency on real hardware. To demonstrate this, we measure the compression time and peak memory consumption of Taco-SVD when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio. These metrics, summarized in Table 3, indicate that Taco-SVD achieves a good balance between compression time and memory consumption. Additionally, to assess the acceleration effect of Taco-SVD, we compared the throughput (tokens generated per second) of original LLaMA2-7B and its compressed version by layer removal and Taco-SVD with different sequence lengths and batch sizes. As shown in Figure 4, Taco-SVD consistently ensures an enhancement in the generation speed and achieves an acceleration effect comparable to pure layer removal. It is important to note that although Taco-SVD retains a small portion of parameters in the redundant layers to mitigate the internal inconsistencies introduced by layer removal, this structure is extremely low-rank, resulting in negligible inference latency while maintaining task performance.
Model
Average
Compression Time(h)
Memory
Consumption(MiB)
LaCO 0.05 45046
ASVD 5.5 17632
SliceGPT 0.6 18195
Taco-SVD 0.16 29050
Table 3: Compression time and peak memory usage of Taco-SVD and baselines for LLaMA2-7B under 25% compression ratio on a single A100 GPU.
Figure 4: Throughput of LLaMA2-7B and Taco-SVD compressed model under 25% compression ratio on a single A100 GPU. Top: Throughput across different sequence lengths (batch size = 32). Bottom: Throughput across different batch sizes (sequence length = 32).
4.3 Ablation Study
Calibration Data
We conducted ablation studies to investigate the impact of calibration dataset and the number of calibration data used for singular value gradient attribution. Table 4 summarizes the results of Taco-SVD when compressing LLaMA3.1-8B-Instruct under 20% compression ratio using different calibration datasets and varying numbers of calibration data from WikiText-2. The results demonstrate that Taco-SVD is robust to both the choice of calibration dataset and the quantity of calibration data. Notably, even with as few as 64 calibration samples, Taco-SVD achieves competitive performance, highlighting its efficiency in leveraging minimal data for effective compression. We report the detailed results in the Appendix A.6.
Ablation Type
Calibration
Dataset
WikiText-2 PTB
Average
Accuracy
Varying Dataset WikiText-2 37.86 63.97 47.12
C4 40.54 71.42 46.17
Varying Number 64 46.5 86.51 47.06
128 39.91 76.41 46.93
256 38.73 79.13 46.67
512 37.86 63.97 47.12
Table 4: Comparison of Taco-SVD using different types and amounts of calibration data for compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
One-shot Pruning vs Iterative Pruning
As detailed in Section 3.4, Taco-SVD employs a layer-wise backward compression strategy to compress redundant layers. This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers. Table 5 shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
Pruning Strategy WikiText-2 PTB
Average
Accuracy
Compression
Time(h)
One-shot Pruning 37.86 63.97 47.12 0.16
Iterative Pruning 38.39 72.18 47.13 0.22
Table 5: Comparison of one-shot pruning and iterative pruning for LLaMA3.1-8B-Instruct under 20% compression ratio.
5 Discussion
5.1 How Taco-SVD Stabilizes Layer Removal and Enhances Model Consistency
Instabilities Caused by Layer Removal
Layer removal can disrupt the internal consistency of a model, leading to performance degradation. In response, the model attempts to compensate for the loss of removed layers, triggering a process we refer to as self-repair (Rushing and Nanda, 2024). While this self-repair mechanism helps the model maintain partial functionality, it is often noisy and incomplete. The process cannot fully recover the removed components, resulting in higher perplexity and reduced cross-task generalization. This phenomenon is task-dependent, with certain tasks being more sensitive to the loss of critical task-relevant transformations.
Taco-SVD as a Stabilizing Solution
Taco-SVD effectively addresses these issues by preserving key task-critical singular value directions, reducing the need for self-repair and stabilizing the model’s internal consistency. Specifically, Taco-SVD:
  • Retains task-critical singular values, minimizing disruption to the residual stream. This helps maintain the original structure and function of the model, reducing perplexity and improving performance across diverse tasks.
  • Mitigates the reliance on self-repair, ensuring that the model does not need to rely on incomplete compensations. By aligning the retained singular values with task-specific objectives, Taco-SVD facilitates smoother adjustments that preserve both task performance and internal consistency.
5.2 Taco-SVD vs LoRA vs LLM-Streamline
Limitations of fine-tuning Pruned LLMs with LoRA
Parameter-efficient fine-tuning methods such as LoRA introduce additional parameters across the remaining layers to compensate for the removed layers. However, training the pruned model with LoRA is a process of redistributing the function of the pruned layers (Chen et al., 2024), these updates often introduce intruder dimensions, which may lead to potential issues such as Spectral Dilution (Shuttleworth et al., 2024): LoRA’s weight updates alter the original singular value spectrum, the new singular vectors added by LoRA are often orthogonal to the original pre-trained singular vectors, which diminishes the contribution of task-agnostic features learned during pre-training.
Limitations of Layer Replacement with Dense Modules
Methods like LLM-Streamline (Chen et al., 2024) train a dense module from scratch to approximate the functionality of redundant layers within LLMs. However, dense modules possess high-dimensional internal structures, which may cause overfitting in scenarios with limited data.
Taco-SVD’s Generalization Advantage via Low-Rank Learning
In contrast, Taco-SVD avoids such issues by preserving the original spectral structure while aligning it with task-specific objectives.
  • Taco-SVD identifies and retains the singular value directions in the redundant layer that are most critical to task performance. It maintains the original pre-trained singular value spectrum, preventing the introduction of orthogonal updates to the original singular vectors in the remaining layers.
  • Previous literature has demonstrated the effectiveness and practicality of low-rank learning in deep networks (Oymak et al., 2019; Kamalakara et al., 2022; Yaras et al., 2024). Therefore, instead of training a dense module from scratch, Taco-SVD fine-tunes only the minimal task-aware subspaces in redundant layers. It can avoid overfitting by constraining the model to learn within a low-rank structure.
6 Conclusion
We propose Taco-SVD, a task-aware framework for compressing large language models by retaining task-critical singular value directions. Taco-SVD mitigates the disruptions caused by traditional layer removal and avoids the need for computationally expensive fine-tuning compensation by preserving key task-critical transformations. Experimental results show that Taco-SVD achieves lower perplexity, better task performance, and improved cross-task robustness with minimal resource overhead. This work highlights the effectiveness of aligning low-rank decompositions with task-specific objectives for efficient and robust model compression.
7 Limitations
While Taco-SVD demonstrates strong performance across various tasks and architectures, we acknowledge certain limitations that highlight opportunities for future research:
Dependence on Calibration Data Taco-SVD relies on calibration data for singular value gradient attribution. In our experiments, using as few as 64 calibration samples from the general-purpose WikiText dataset achieved strong performance across multiple tasks. However, while this approach works well in these settings, we have not yet explored whether using domain-specific calibration data could yield better results in specialized domains. Further testing with such data may provide valuable insights into its potential advantages.
Trade-Off Between Compression and Retained Parameters To ensure robustness and task consistency, Taco-SVD retains a small fraction of low-rank parameters in redundant layers, mitigating disruptions caused by pure layer removal. This design choice introduces a slight trade-off in terms of parameter retention, which could limit its suitability for applications requiring extreme compression ratios. Further reduction techniques may complement Taco-SVD in such scenarios.
Scalability to Extremely Large Models While Taco-SVD performs effectively on models like LLaMA2 and LLaMA3, its application to ultra-large-scale models (e.g., hundreds of billions of parameters) could pose additional computational challenges. Future work could focus on optimizing the algorithm for such architectures, enabling its application to next-generation LLMs.
Ethical Considerations
Our research adheres to the ACL Code of Ethics, ensuring transparency, responsible use of data, and consideration of potential social impacts. All datasets used in this work are publicly available and have been appropriately cited, ensuring compliance with data usage agreements and privacy regulations.
While Taco-SVD is designed to optimize the efficiency and scalability of large language models, we recognize that such technologies could be misused in applications that may perpetuate harmful biases or deploy models in contexts lacking adequate oversight. To mitigate these risks, we advocate for responsible deployment practices, including thorough testing and monitoring for unintended biases.
Moreover, we acknowledge the computational resources required for training and testing large language models. To minimize environmental impact, we conducted experiments on energy-efficient hardware (NVIDIA A100 GPUs) and report our computational cost transparently. Further details can be found in the Appendix.
Appendix A Appendix
A.1 The gradient of singular values
For a weight matrix $W\in\mathbb{R}^{m\times n}$ in the selected redundant layers, its differential form can be expressed as:
$\partial W=\partial U\Sigma V^{T}+U\partial\Sigma V^{T}+U\Sigma\partial V^{T}$
$U^{T}\partial WV=U^{T}\partial U\Sigma+\partial\Sigma+\Sigma V^{T}\partial V$
Since both $U$ and $V$ are orthogonal matrices, we have:
$U^{T}U=I_{m},\quad V^{T}V=I_{n}$
$\partial U^{T}U+U^{T}\partial U=O_{m},\quad\partial V^{T}V+V^{T}\partial V=O_{n}$
This implies that $U^{T}dU$ and $dV^{T}V$ are asymmetric matrices. Therefore, the diagonal elements of $U^{T}dU\Sigma$ and $\Sigma V^{T}dV$ are zero, leading to the diagonal elements of $U^{T}\partial WV$ being:
$I_{k}\odot U^{T}\partial WV=\partial\Sigma$
where $I_{k}$ represents the $k\times k$ identity matrix, $\odot$ denotes element-wise multiplication.
For a singular value $\sigma_{i}$, its differential form can be written as:
$\partial{\sigma_{i}}=u_{i}^{T}\partial Wv_{i}$
Since $\sigma_{i}$ is a scalar, we have:
$\partial\sigma_{i}$ $=\text{tr}(\partial\sigma_{i})$
$=\text{tr}(u_{i}^{T}\partial Wv_{i})$
$=\text{tr}[(u_{i}v_{i}^{T})^{T}\partial W]$
thereby, the derivative of $\sigma_{i}$ with respect to W is:
$\frac{\partial\sigma_{i}}{\partial W}=u_{i}v_{i}^{T}$
For a calibration dataset $D$, the gradient of a singular value $\sigma_{i}$ with respect to the task loss can be interpreted as the projection of the weight gradient matrix $G$ onto the corresponding singular direction, given by:
$\frac{\partial L}{\partial\sigma_{i}}=u_{i}^{T}\frac{\partial L}{\partial W}v_{i}$
Then, for all the singular values $\Sigma$, we have:
$\frac{\partial L}{\partial\Sigma}=I_{k}\odot U^{T}\left(\frac{\partial L}{\partial W}\right)V$
A.2 Layer-wise Low-Rank Decomposition
This section provides the implementation of Layer-wise Low-Rank Decomposition, as detailed in Algorithm 1.
Algorithm 1 Task-Aware SVD for LLM Compression
1:Model $M$ with $N$ layers, Calibration Dataset $D$, Retain Ratio $r$
2:$\text{W}\in\{W_{q},W_{k},W_{v},W_{o},W_{up},W_{down}\}$ for weight matrices in self-attention and MLP blocks
3:Compressed Model $\tilde{M}$
4:Step 1: Redundant Layer Selection
5:for $i=1$ to $N$ do
6:     $\text{cos}_{i}\leftarrow\frac{H_{i}^{T}H_{i+1}}{\|H_{i}\|_{2}\|H_{i+1}\|_{2}}$
7:end for
8:Select $L$ layers with highest cosine similarity.
9:Step 2: Task-aware SVD Compression:
10:for each layer $l\in L$, starting from the last layer do
11:     $\text{W}\in\{W_{q},W_{k},W_{v},W_{o},W_{up},W_{down}\}$
12:     $\text{W}\leftarrow\mathbf{U}\Sigma\mathbf{V^{T}}$
13:     $\text{G}\leftarrow\frac{\partial\mathcal{L}_{\text{}}}{\partial W}$   # compute gradient matrices using $D$
14:     $\frac{\partial\mathcal{L_{\text{}}}}{\partial\sigma_{i}}\leftarrow\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$   # Compute gradients for singular values
15:     $\tilde{W}\leftarrow\tilde{U}\tilde{\Sigma}\tilde{V^{T}}$   # Truncate singular values based on Eq.5, retain only $r\%$ parameters per matrix
16:end for
17:return $\tilde{M}$
Figure 5: Validation loss curves of different methods on the Alpaca dataset (1 epoch)
A.3 Experimental Setup and Hyperparameters Configuration
To ensure a fair comparison, all experimental setups are consistent across all methods. Below we describe the experimental setup and hyperparameters configuration in detail.
Training-free Setting
We report the results of Taco-SVD and other training-free baselines under the compression ratio of 20% in Table 1. Specifically, we selected the same 256 samples from the WikiText-2 dataset as calibration data.
Training-based Setting
Table 6 provides the detailed configurations of training-based methods under the compression ratio of 20%. All our experiments are conducted on a single NVIDIA A100 GPU with mixed precision enabled.
HyperParameters Setting
Dataset Alpaca
Huggingface Dataset Path yahma/alpaca-cleaned
Batch Size 32
Micro Batch Size 4
Epochs 1
Learning Rate 3.00E-04
Max Length 256
Train on Inputs TRUE
Add EOS Token FALSE
LoRA-Rank 256
LoRA-Alpha 16
LoRA-Dropout 0.05
LoRA-Target-Modules
q_proj, k_proj, v_proj, o_proj,
up_proj, down_proj, gate_proj
Prompt-Template Alpaca Template
Table 6: Experimental setup and hyperparameter configurations.
Method Ratio WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Dense 0% 5.47 20.83 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
LaCo 18% 39.43 93.98 26.2 48.36 59.19 41.93 32.42 68.72 24.02 42.98
SliceGPT 21.50% 8.63 99.25 22 54.34 60.62 37.4 27.99 63.38 23.18 41.27
ASVD 20% 9.7 88.39 27.8 58.41 61.09 46.97 32.42 71.05 25.09 46.12
Taco-SVD 21.60% 16.12 44.09 24 53.59 62.9 42.94 32.76 71.06 22.88 44.30
LaCo 24% 50.33 102.84 25 45.33 60.06 38.3 29.69 65.51 23.99 41.13
SliceGPT 25.10% 9.54 118.17 20.6 50.91 59.63 35.66 26.41 61.62 23.43 39.75
ASVD 25% 53.48 374.45 17 39.35 53.12 32.64 22.78 61.48 22.95 35.62
Taco-SVD 25% 30.2 60.74 21.6 46.59 61.8 40.41 28.84 66.1 23.18 41.22
LaCo 30% 94.5 122.09 22 42.26 59.43 35.7 28.41 62.89 24.32 39.29
SliceGPT 29.10% 10.91 145.24 18.8 47.56 58.64 34.18 25.34 59.58 23.15 38.18
ASVD 30% 1877.63 3340.65 14.4 27.78 50.59 26.37 21.84 51.58 20.88 30.49
Taco-SVD 30% 35.38 88.96 19.4 42.72 61.17 38.43 27.73 63.98 23.18 39.52
LaCo 36% 138.82 164.02 19 38.34 58.96 33.71 27.13 60.55 23.55 37.32
SliceGPT 33.20% 12.8 176.54 17.8 44.23 57.61 32.28 23.55 58.59 22.68 36.68
ASVD 35% 2599.79 6529.72 13.6 26.81 50.83 25.94 23.12 53.59 19.33 30.46
Taco-SVD 35.00% 68.43 125.79 18.8 40.25 60.35 36.77 27.18 61.11 22.63 38.16
LaCo 39% 386.55 289.98 18.4 33.96 56.99 31.34 26.88 58.6 23.45 35.66
SliceGPT 41.40% 19.1 257.66 15.8 35.94 53.43 29.62 20.82 55.98 22.01 33.37
ASVD 40% NaN NaN 13.4 25.76 51.3 26.1 21.67 54.19 20.64 30.44
Taco-SVD 40.10% 80.42 247.55 18 36.83 59.59 33.31 26.62 59.63 21.88 36.55
Table 7: Zero-shot performance of LLaMA2-7B compressed by Taco-SVD and training-free baselines under 20% to 40% compression. Bold values indicate the best performance.
A.4 Validation Loss Curves
We plot the validation loss curves when training pruned Mistral-7B with 20% compression ratio with different compensation methods. As shown in Figure 5, we can observe that Taco-SVD has better initialization, leading to more stable and faster convergence during fine-tuning.
A.5 Results under High Compression Ratio
In this section, we present the results of Taco-SVD compared with the training-free baselines using LLaMA2-7B under high compression ratios, as shown in Table 7. The results demonstrate that our proposed method consistently outperforms the baselines, even at high compression ratios of up to 40%.
A.6 Robustness of Taco-SVD towards different Calibration Dataset
In this section, we provide details of the ablation studies conducted to investigate the impact of calibration datasets and the number of calibration data used for singular value gradient attribution. Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of Taco-SVD when compressing LLaMA3.1-8B-Instruct under 20% compression ratio. Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of Taco-SVD to the change in the number of calibration data. All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage. As shown in Figure 6, we can observe that Taco-SVD consistently achieves strong performance, indicating that our method is robust to variations in both the calibration dataset and the number of calibration data. Tables 8 and 9 summarize the results of Taco-SVD when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
Figure 6: Performance of Taco-SVD on LLaMA3.1-8B-Instruct under 20% compression using (a) different calibration datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2. Taco-SVD demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
WikiText-2 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
C4 40.54 71.42 24 57.91 67.72 42.11 36.69 67.25 27.5 46.17
Table 8: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by Taco-SVD under 20% compression using 512 samples from WikiText-2 and C4 as calibration datasets.
Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
WikiText-2-64 46.5 86.51 22.6 59.97 69.3 44.36 37.29 68.5 27.37 47.06
WikiText-2-128 39.91 76.41 22.8 60.23 69.93 44.24 35.92 67.9 27.5 46.93
WikiText-2-256 38.73 79.13 21.8 59.89 70.24 44.23 36.26 67.19 27.07 46.67
WikiText-2-512 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
Table 9: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by Taco-SVD under 20% compression with varying calibration data sizes (64, 128, 256, 512) from WikiText-2.

### v2

@@
-Rethinking Layer Removal: Preserving Critical Components with Task-Aware Singular Value Decomposition
+Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression
 Abstract
-Layer removal has emerged as a promising approach for compressing large language models (LLMs) by leveraging redundancy within layers to reduce model size and accelerate inference.
-However, this technique often compromises internal consistency, leading to performance degradation and instability, with varying impacts across different model architectures.
-In this work, we propose Taco-SVD, a task-aware framework that retains task-critical singular value directions, preserving internal consistency while enabling efficient compression.
-Unlike direct layer removal, Taco-SVD preserves task-critical transformations to mitigate performance degradation.
-By leveraging gradient-based attribution methods, Taco-SVD aligns singular values with downstream task objectives.
-Extensive evaluations demonstrate that Taco-SVD outperforms existing methods in perplexity and task performance across different architectures while ensuring minimal computational overhead.
+Layer removal is an effective technique for compressing large language models (LLMs) by reducing redundancy and improving inference efficiency.
+However, indiscriminate pruning disrupts representation stability, leading to performance degradation.
+We propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), which preserves representation-critical singular values to mitigate these effects.
+Unlike direct layer removal, GRASP leverages gradient-based attribution on a syntax- and semantics-rich dataset to guide the selection of representation-critical singular values.
+By selectively applying singular value decomposition (SVD) to affected layers, GRASP achieves efficient compression while maintaining representation stability with minimal overhead.
+Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods in perplexity and downstream task performance.
+The source code is available at GRASP.
 […]
 […]
-Rethinking Layer Removal: Preserving Critical Components with Task-Aware Singular Value Decomposition
+Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression
 Kainan Liu1,2,†, Yong Zhang1,†, Ning Cheng1,*, Zhitao Li1, Shaojun Wang1, Jing Xiao1, 1Ping An Technology (Shenzhen) Co., Ltd., China 2The Hong Kong University of Science and Technology (Guangzhou) {zhangyong203, chengning211}@pingan.com.cn
@@
 1 Introduction
-Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing tasks, including language generation, reasoning, and question answering (Brown et al., 2020; Touvron et al., 2023).
-However, their massive parameter sizes create significant computational and memory challenges, making them difficult to deploy on resource-constrained devices (Zhou et al., 2024).
+Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, including language generation, reasoning, and question answering (Brown et al., 2020; Touvron et al., 2023b).
+However, their massive parameter sizes pose computational and memory challenges, hindering deployment on resource-limited devices (Zhou et al., 2024).
 To address this, model compression techniques such as quantization (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023), knowledge distillation (Gu et al., 2023; Xu et al., 2024), and pruning(Sun et al., 2024; Ashkboos et al., 2024) have been widely explored.
 Among these, structured pruning methods, especially layer removal (Men et al., 2024; Yang et al., 2024; Song et al., 2024), have shown promise by leveraging redundancy within layers of LLMs to achieve significant parameter reductions and inference speedup.
-Figure 1: Zero-shot performance of Taco-SVD and pure layer removal under 20% compression ratio.
+Figure 1: Zero-shot performance of GRASP and pure layer removal under 20% compression ratio.
 Dashed lines represent perplexity on WikiText-2, while bar charts show the percentage of the original model’s average accuracy on seven reasoning datasets.
-Despite its simplicity and computational efficiency, pure layer removal introduces two notable challenges, as shown in Figure 1.
-First, while task accuracy may remain stable in certain cases, perplexity (PPL) often degrades significantly, reflecting disruptions to the model’s internal consistency, which stems from impaired information flow and misaligned intermediate representations (Liang et al., 2024).
-This increase in PPL suggests that layer removal may affect critical transformations responsible for maintaining coherent representations, potentially reducing the model’s generalization ability on complex or unseen tasks.
-Second, the impact of layer removal varies significantly across model architectures.
-For instance, under the same pruning ratio, LLaMA2 (Touvron et al., 2023) maintains reasonable performance, whereas LLaMA3 (Dubey et al., 2024) experiences a steep performance drop.
-This observation indicates that the tolerance to layer pruning is architecture-dependent.
-Such variability underscores a lack of robustness in pure layer removal, where certain task-critical transformations may be unevenly affected, resulting in performance inconsistencies across different architectures.
-Our analysis suggests that these issues arise from the removal of transformations that play an essential role in maintaining both internal consistency and task-specific performance.
-While existing methods, such as LoRA-based fine-tuning (Hu et al., 2022) or lightweight layer replacements (Chen et al., 2024), aim to recover or approximate the lost functionality, they often incur substantial computational costs and do not fully address the challenge of preserving task-critical transformations in a resource-efficient manner.
-To address these challenges, we propose Taco-SVD (Task-Aware Singular Value Decomposition for LLM Compression), a novel framework designed to address the limitations of pure layer removal by preserving task-critical transformations.
-Taco-SVD fundamentally rethinks layer removal by integrating task-aware optimization into singular value decomposition, ensuring that compression decisions align with task-specific objectives.
-Unlike conventional SVD, which focuses on global matrix reconstruction, Taco-SVD leverages gradient attribution to directly quantify the task-specific importance of singular values.
-Using a small calibration dataset from a general-purpose, unlabeled source, our method identifies and preserves the most impactful transformations within redundant layers.
-This approach minimizes the indiscriminate loss of critical task-specific transformations, maintaining internal consistency and representation flow while achieving efficient compression.
-To validate our approach, we conducted experiments on tasks such as language generation, commonsense reasoning, and mathematical reasoning using LLaMA and Mistral models.
-Taco-SVD achieves parameter reduction comparable to pure layer removal while improving perplexity and task performance.
-Notably, Taco-SVD demonstrates consistent robustness across diverse model architectures, maintaining strong performance without additional training.
-It also achieves strong performance with minimal calibration data—only 64 samples from a general-purpose, unlabeled dataset, demonstrating its effectiveness.
-Moreover, fine-tuning further enhances its performance, highlighting Taco-SVD’s robustness and efficiency in compressing large language models.
-Figure 2: Overview of Taco-SVD.
-The figure illustrates how Taco-SVD uses gradient attribution to identify task-critical singular values and applies SVD to compress redundant layers while preserving task-relevant transformations.
+Despite its simplicity and efficiency, layer removal introduces two main challenges, as shown in Figure 1.
+First, while task accuracy may remain stable in some cases, the perplexity (PPL) often degrades significantly, reflecting disruptions in the model’s internal representation consistency.
+This degradation arises from impaired information flow and misaligned intermediate representations (Liang et al., 2024).
+The sharp increase in PPL suggests that layer removal eliminates critical transformations necessary for maintaining coherent representation structures, potentially limiting the model’s generalization on complex or unseen tasks.
+Second, the impact of layer removal varies significantly across models.
+For example, under the same pruning ratio, LLaMA2 (Touvron et al., 2023b) maintains reasonable performance, while LLaMA3 (Dubey et al., 2024) experiences a steep performance drop.
+This discrepancy suggests that layer removal affects models unevenly, as certain structural transformations are more essential for maintaining stable representations, leading to performance inconsistencies across models.
+Our analysis suggests that these issues arise from the removal of transformations that play an essential role in preserving internal consistency and maintaining structured representations.
+In this paper, we introduce GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel hybrid pruning framework that integrates layer removal with gradient-based singular value selection for efficient LLM compression.
+Unlike conventional pruning techniques that either remove full layers or apply global SVD compression, GRASP selectively applies singular value decomposition (SVD) only to layers affected by pruning.
+A key component of GRASP is its use of a small, syntax- and semantics-rich dataset (e.g., WikiText-2), which serves as a representation evaluation dataset to guide the selection of singular values.
+Instead of relying solely on magnitude-based ranking, GRASP leverages gradient attribution computed from this dataset to dynamically identify and retain singular values critical for preserving language representations.
+This ensures that despite layer removal, the model retains essential syntactic and semantic structures, minimizing degradation in perplexity and downstream task performance.
+We evaluate GRASP on 10 datasets and five models from two distinct LLM families (LLaMA, Mistral).
+Extensive experiments validate the effectiveness of GRASP.
+It achieves compression rates comparable to pure layer removal while significantly improving perplexity and task performance across a wide rage of downstream tasks.
+Notably, GRASP maintains strong robustness across different models without requiring additional training.
+Furthermore, it achieves strong performance with minimal calibration data—using only 64 samples from a general-purpose, unlabeled dataset, demonstrating the efficiency of our approach.
+Additionally, post-training compensation further enhances GRASP’s performance, highlighting its adaptability in compressing large language models while preserving their linguistic capabilities.
+Figure 2: (A): Overview of GRASP.
+The figure illustrates how GRASP uses gradient attribution to identify representation-critical singular values and applies SVD to compress redundant layers while preserving representation-relevant transformations.
+(B): SVD-based LLM compression and (C): Layer removal compression methods
 2 Related Work
-2.1 Layer Removal Techniques and Their Impact on Model Consistency
+2.1 Layer Removal-based LLM Compression
 Layer removal is a structured pruning technique that eliminates redundant layers within Large Language Models (LLMs), aiming to reduce the model’s computational and memory requirements.
@@
 LaCo (Yang et al., 2024), on the other hand, proposes a gradual compression approach, progressively merging redundant layers from deeper to shallower parts of the network.
-Although effective in reducing model size, layer removal often disrupts internal consistency, leading to degraded task performance and increased perplexity, as shown in our analysis building on the concept introduced by Liang et al.
+Although effective in reducing model size, layer removal disrupts representation coherence, leading to performance degradation and increased perplexity, as analyzed by Liang et al.
 (2024).
-Additionally, its sensitivity to different architectures limits generalizability, with some models tolerating pruning better than others.
-These challenges highlight the need for approaches that preserve task-critical information while ensuring robust compression across diverse scenarios.
-2.2 Post-Pruning Compensation and Training-Based Recovery Approaches
-To mitigate the degradation caused by layer removal, post-pruning compensation methods have been developed.
-(Kim et al., 2024) introduced Shortened LLaMA, which incorporates LoRA (Hu et al., 2022) to recover the ability of layer removal models.
-However, LoRA alters the original singular value spectrum, which may diminish the contribution of pre-trained features (Shuttleworth et al., 2024).
-To address this issue, LLM-Streamline (Chen et al., 2024) trains a lightweight module such as feed-forward neural network (FFN) or a transformer layer to approximate the functionality of the pruned redundant layers.
-Although effective to some extent, these methods incur significant computational and data costs, limiting their practicality in resource-constrained scenarios.
-2.3 Singular Value Decomposition for LLM Compression
-Singular Value Decomposition (SVD) is a foundational technique for compressing dense models by decomposing weight matrices into low-rank structures.
-However, vanilla SVD focuses solely on global matrix reconstruction, providing general approximations without considering task-specific characteristics.
-Recent methods have enhanced SVD to better align with model and task requirements.
+To mitigate this, post-training compensation methods have been proposed.
+Kim et al., 2024 introduced Shortened LLaMA, which employs LoRA (Hu et al., 2022) to restore pruned models’ capabilities.
+However, LoRA modifies the singular value spectrum, potentially weakening pre-trained features (Shuttleworth et al., 2024).
+LLM-Streamline (Chen et al., 2024) addresses this by training a lightweight module, such as an FFN or transformer layer, to approximate the removed layers.
+While effective, these methods impose high computational and data costs, limiting feasibility in resource-constrained settings.
+2.2 SVD-based LLM Compression
+Singular Value Decomposition (SVD) is widely used in model compression to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss.
+Recent methods have enhanced SVD to reduce LLM compression error.
 FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD.
-ASVD (Yuan et al., 2023) uses activation patterns from a calibration dataset to scale weight matrices, reducing compression-induced activation errors.
+ASVD (Yuan et al., 2023) uses activation patterns from a representation evaluation dataset to scale weight matrices, reducing compression-induced activation errors.
 SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss.
 Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices.
-Despite these advancements, existing methods mainly focus on optimizing compression loss without directly addressing the relationship between singular values and task-specific losses.
-Taco-SVD bridges this gap by integrating task-aware optimization into SVD, leveraging gradient attribution to retain task-critical singular values and align compressed models with task objectives.
+GRASP differs from these methods in three key aspects.
+First, rather than minimizing F2 norm reconstruction loss, GRASP optimizes a representation learning attribution loss, computed on a small representation evaluation dataset to ensure selection aligns with model representation rather than weight approximation.
+Second, instead of using a top-k singular value selection strategy, GRASP leverages gradient attribution to dynamically retain representation-critical singular values, which are not necessarily the largest in magnitude.
+Third, while SVD-based pruning is typically applied to all layers, GRASP is selectively used only in pruned layers, ensuring that layer removal does not disrupt critical transformations.
 3 Method
 3.1 Overview and Motivation
-In this paper, we propose a novel task-aware compression framework named Taco-SVD, designed to compress redundant layers in large language models (LLMs) while preserving task-critical transformations.
-Unlike traditional layer removal methods that directly eliminate redundant components, Taco-SVD integrates task-aware singular value decomposition into the layer removal paradigm.
-By introducing gradient-based attributions, it systematically aligns the importance of singular values with task-specific objective, ensuring that the most essential components within redundant layers are retained during truncation.
-Figure 2 provides an overview of Taco-SVD.
-The method consists of two main steps: (1) identifying redundant layers based on cosine similarity, and (2) applying task-aware singular value decomposition (SVD) to retain the most critical singular values and their corresponding singular vectors.
+In this paper, we propose GRASP, a novel hybrid gradient-guided compression framework designed to compress redundant layers in LLMs while preserving representation-critical transformations.
+Figure 2 provides an overview of GRASP.
+The method consists of two main steps: (1) identifying redundant layers based on cosine similarity, and (2) selecting and retaining representation-critical singular values and their corresponding singular vectors using gradient-based attribution.
 Below, we describe each step in detail.
 3.2 Redundant Layer Selection
-The first step in Taco-SVD is identifying redundant layers.
+The first step in GRASP is identifying redundant layers.
 These are layers that contribute minimally to the transformation of hidden states, exhibiting high redundancy and limited impact on overall model performance.
@@
 A high cosine similarity indicates minimal transformation, suggesting that the layer is redundant.
-Instead of directly removing these layers, Taco-SVD compresses their weight matrices using a task-aware approach to preserve task-relevant transformations.
-3.3 Task-Aware Singular Value Decomposition
-3.3.1 Motivation for Task-Aware SVD
-Singular value decomposition (SVD) is a standard technique for compressing weight matrices by providing a global low-rank approximation.
-Traditional SVD methods retain the top-$k$ singular values based solely on their magnitudes, which are task-agnostic and often fail to capture task-relevant components.
-This can lead to the removal of critical transformations necessary for downstream tasks (Hsu et al., 2022; Wang et al., 2024).
-To address this limitation, Taco-SVD introduces task awareness into SVD by using gradient-based attribution to evaluate the task-specific importance of each singular value and its corresponding singular vectors.
-This ensures that only the most task-relevant transformations are preserved.
+Instead of directly removing these layers, GRASP compresses weight matrices with a gradient-based approach to retain critical internal transformations.
+3.3 Gradient-Guided Singular Value Selection
+3.3.1 Representation Learning Attribution Loss (RLAL)
+To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value based on its contribution to model performance, rather than relying on magnitude alone.
+Unlike traditional SVD-based methods that select top-$k$ singular values using a reconstruction loss measured by the Frobenius norm ($F_{2}$ norm), GRASP utilizes a Representation Learning Attribution Loss (RLAL), designed specifically for general representation capability modeling.
+RLAL is computed on a small, unlabeled representation evaluation dataset derived from corpora rich in syntactic and semantic structures, such as WikiText-2 and similar datasets.
+These datasets ensure that the retained singular values are those most crucial for maintaining the model’s ability to capture fundamental language structures, including syntax and semantics, even after compression.
+Unlike task-specific objectives, RLAL focuses on preserving linguistic consistency and general-purpose representation learning rather than optimizing for any particular downstream application.
+Mathematically, RLAL is formulated as:
+$\mathcal{L}_{\text{RLAL}}=-\sum_{t}\log P(y_{t}\mid x_{\leq t})$ (2)
+where $y_{t}$ is the target token at time step $t$, $x_{\leq t}$ denotes the sequence of preceding tokens, and $P(y_{t}\mid x_{\leq t})$ represents the predicted probability of the next token given the context.
+This loss function aligns with the standard language modeling loss, ensuring that the model learns meaningful representations.
 3.3.2 Singular Value Decomposition (SVD)
-For a redundant layer’s weight matrix $W\in\mathbb{R}^{m\times n}$, SVD decomposes $W$ as:
-$W=U\Sigma V^{T}$ (2)
+For a redundant layer’s weight matrix $W_{\text{redundant}}\in\mathbb{R}^{m\times n}$, SVD decomposes $W_{\text{redundant}}$ as:
+$W_{\text{redundant}}=U\Sigma V^{T}$ (3)
 where $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{k})$ contains the singular values ($\sigma_{i}\geq 0$), and $U\in\mathbb{R}^{m\times m}$, $V\in\mathbb{R}^{n\times n}$ are orthogonal matrices containing the left and right singular vectors, respectively.
-3.3.3 Task-Aware Singular Value Importance
-To assess the task-specific importance of each singular value, we introduce $\Phi(\sigma_{i})$, the Task-Aware Singular Value Importance, which quantifies the contribution of $\sigma_{i}$ to the task loss $\mathcal{L}_{\text{task}}$.
-Using its first-order Taylor expansion, the alteration in the task loss $L$ caused by $\sigma_{i}$ is approximated as:
-$\Delta\mathcal{L}(\sigma_{i})\approx\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}\Delta\sigma_{i}$ (3)
-where $\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}$ represents the gradient sensitivity of $\sigma_{i}$ to the task loss, and $\Delta\sigma_{i}$ is the perturbation applied to $\sigma_{i}$.
-The gradient sensitivity is calculated by projecting the gradient of $\mathcal{L}_{\text{task}}$ with respect to $W$ onto the singular directions:
-$\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}=\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$ (4)
-where $G=\frac{\partial\mathcal{L}_{\text{task}}}{\partial W}$ is the gradient of the task loss with respect to $W$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the left and right singular vectors corresponding to $\sigma_{i}$.
+3.3.3 Representation-Critical Singular Value Importance
+To assess the representation-critical importance of each singular value, we introduce $\Phi(\sigma_{i})$, the Representation-Critical Singular Value Importance, which quantifies the contribution of $\sigma_{i}$ to the $\mathcal{L}_{\text{RLAL}}$.
+Using its first-order Taylor expansion, the alteration in the $\mathcal{L}_{\text{RLAL}}$ caused by $\sigma_{i}$ is approximated as:
+$\Delta\mathcal{L}_{\text{RLAL}}(\sigma_{i})\approx\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}\Delta\sigma_{i}$ (4)
+where $\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}$ represents the gradient sensitivity of $\sigma_{i}$ to the representation learning attribution loss , and $\Delta\sigma_{i}$ is the perturbation applied to $\sigma_{i}$.
+The gradient sensitivity is calculated by projecting the gradient of $\mathcal{L}_{\text{RLAL}}$ with respect to $W_{\text{redundant}}$ onto the singular directions:
+$\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}=\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$ (5)
+where $G=\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial W}$ is the gradient of the $\mathcal{L}_{\text{RLAL}}$ with respect to $W_{\text{redundant}}$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the left and right singular vectors corresponding to $\sigma_{i}$.
 The detailed proof can be found in the Appendix A.1.
@@
 $\Phi(\sigma_{i})=\sigma_{i}\cdot\left(\mathbf{u}_{i}^{T}G\mathbf{v}_{i}\right)$ (6)
-where $\Phi(\sigma_{i})$ measures the overall contribution of $\sigma_{i}$ to task performance, considering both its magnitude and alignment with the task gradient.
-Larger value of $\Phi(\sigma_{i})$ indicates greater importance.
-We rank singular values in descending order of task relevance based on $\Phi(\sigma_{i})$.
+where $\Phi(\sigma_{i})$ measures the overall contribution of $\sigma_{i}$ to representation performance, considering both its magnitude and alignment with the gradient of the general-purpose model loss.
+A larger value of $\Phi(\sigma_{i})$ indicates greater importance.
+We rank singular values in descending order of representation relevance based on $\Phi(\sigma_{i})$.
 Only the top-$k$ singular values with the highest $\Phi(\sigma_{i})$ are retained, and the compressed weight matrix is reconstructed as:
-$W_{\text{task}}=\sum_{i=1}^{k}\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{T}$ (6)
-This ensures that $W_{\text{task}}$ preserves the transformations most critical to task performance, achieving efficient compression while maintaining task consistency.
+$W_{\text{RLAL}}=\sum_{i=1}^{k}\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{T}$ (7)
+This ensures that $W_{\text{RLAL}}$ preserves the transformations most critical to representation stability, achieving efficient compression while maintaining representation consistency.
 3.4 Layer-wise Low-Rank Decomposition
-Taco-SVD applies a layer-wise backward compression strategy, starting from the last redundant layer and moving sequentially backward.
-For each redundant layer, weight matrices are compressed using Task-Aware Singular Value Decomposition, retaining only the top-$k$ singular values ranked by $\Phi(\sigma_{i})$.
-This ensures a low-rank approximation that preserves task-relevant transformations.
+GRASP applies a layer-wise backward compression strategy, starting from the last redundant layer and moving sequentially backward.
+For each redundant layer, weight matrices are compressed using Gradient-Guided Singular Value Decomposition, retaining only the top-$k$ singular values ranked by $\Phi(\sigma_{i})$.
+This ensures a low-rank approximation that preserves representation-critical transformations.
 By processing one layer at a time, this approach significantly reduces memory overhead, avoiding the need to load the entire model into memory.
@@
 4 Experiments
+In this section, we conduct comprehensive experiments to evaluate GRASP from four key perspectives.
+(1) We first compare our method with several popular pruning-based LLM compression methods to demonstrate its effectiveness (Section 4.2).
+(2) Next, we compare GRASP with SVD-based LLM compression methods under different compression ratios with different LLMs (Section 4.3).
+(3) Then, we further evaluate GRASP on the LongBench (Bai et al., 2023) datasets to assess its performance on long-form text processing tasks (Section 4.4).
+(4) Finally, we analyze the inference speed-up achieved by GRASP (Section 4.5).
 4.1 Experimental Setup
-To validate the effectiveness of Taco-SVD, we conducted experiments across multiple large language models (LLMs) and datasets, focusing on evaluating task performance, cross-model robustness, and computational efficiency.
-Below, we detail the models, datasets, baselines, and experimental configurations.
+Below we detail the models, benchmarks and implementation details, with more experimental setups provided in Appendix A.3 due to page limit.
 Models
-We evaluated our method on a diverse set of LLMs including LLaMA family models LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023), LLaMA3.1-8B-Instruct (Dubey et al., 2024) and Mistral-7b (Jiang et al., 2023).
-These models enable a comprehensive assessment of our method across varying sizes and architectures, highlighting its generability and robustness.
-Datasets
-Ten datasets are used as evaluation benchmarks including three language generation datasets: WikiText-2 (Merity et al., 2017), PTB (Marcus et al., 1993), C4 (Raffel et al., 2020) and seven datasets for common sense and mathematical reasoning include: WinoGrande (Sakaguchi et al., 2020), Hellaswag (Zellers et al., 2019), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-c, ARC-e (Clark et al., 2018) for common sense reasoning, and MathQA (Amini et al., 2019) for mathematical reasoning.
-Metric
-We use perplexity and accuracy for language generation tasks and reasoning tasks, respectively.
-All reasoning tasks are tested in zero-shot setting using the LM-Evaluation-Harness framework (Gao et al., 2024).
-Baselines
-We compare Taco-SVD against training-free and training-based compression methods at the same compression ratio.
-For training-free methods, we select the layer removal technique LaCo (Yang et al., 2024), as well as a structured pruning method SliceGPT (Ashkboos et al., 2024) and the SVD-based LLM compression method ASVD (Yuan et al., 2023).
-For training-based methods, we compare Taco-SVD with LoRA (Hu et al., 2022) and the layer replacement method LLM-Streamline (Chen et al., 2024) with two variants: LLM-Streamline-FFN and LLM-Streamline-Layer, which replace redundant layers with Feed-Forward Neural Network (FFN) and Transformer Layer (Layer), respectively.
+We evaluated our method on a diverse set of LLMs including LLaMA family models LLaMA-7B (Touvron et al., 2023a), LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023b), LLaMA3.1-8B-Instruct (Dubey et al., 2024) and Mistral family models Mistral-7B (Jiang et al., 2023).
+Benchmark
+10 datasets are used as evaluation benchmarks including two pre-training datasets: WikiText-2 (Merity et al., 2017), PTB (Marcus et al., 1993), seven common sense reasoning datasets include: WinoGrande (Sakaguchi et al., 2020), Hellaswag (Zellers et al., 2019), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-c, ARC-e (Clark et al., 2018), MathQA (Amini et al., 2019) and one generation benchmark GSM8K (Cobbe et al., 2021).
+Additionally, we include LongBench (Bai et al., 2023) to assess LLM’s performance on long-form text processing after pruning.
 Implementation Details
-To ensure a fair comparison, all experimental setups are consistent across all methods.
-Specifically, for training-free methods, we randomly select 512 samples from the WikiText-2 dataset as calibration data.
-For training-based methods, all models are fine-tuned on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32.
-We use AdamW (Loshchilov and Hutter, 2019) as our optimizer and set the learning rate to $3\times 10^{-4}$.
-All our experiments are conducted on a single NVIDIA A100 GPU with mixed precision enabled.
-Further experimental details can be found in the Appendix A.3.
-Model Method Ratio WikiText-2 PTB C4 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
-Mistral-7B Dense 0% 5.25 23.73 8.36 32.6 80.85 73.88 61.22 50.43 80.58 35.88 59.35
-LaCo 21.1% 244.53 524.26 243.98 20.2 35.17 57.54 26.23 24.66 52.77 24.29 34.41
-SliceGPT 20% 9.04 107.65 38.17 19.2 51.05 59.43 34.84 25.26 60.72 23.25 39.11
-ASVD 20% 19.28 171.54 27.93 19.8 55.68 55.96 35.84 24.49 66.21 24.19 40.31
-Taco-SVD 20.3% 26.42 77.98 28.88 21.4 51.47 68.19 42.82 35.24 66.59 25.83 44.51
-LLaMA2-7B Dense 0% 5.47 20.83 7.29 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
-LaCo 18.1% 39.43 93.98 71.77 26.2 48.36 59.19 41.93 32.42 68.72 24.02 42.98
-SliceGPT 21.5% 8.63 99.25 40.7 22 54.34 60.62 37.4 27.99 63.38 23.18 41.27
-ASVD 20% 9.7 88.39 12.46 27.8 58.41 61.09 46.97 32.42 71.05 25.09 46.12
-Taco-SVD 21.6% 16.12 44.09 23.13 24 53.59 62.9 42.94 32.76 71.06 22.88 44.30
-LLaMA-2-13B Dense 0% 4.88 28.92 6.99 32.4 73.27 69.61 59.72 45.56 78.78 30.02 55.62
-LaCo 19.5% 27.26 59.14 27.42 27.6 52.25 62.98 42.52 33.36 69.91 25.32 44.85
-SliceGPT 21.5% 7.44 99.77 37.65 29.4 58.92 64.8 38.96 31.74 64.42 23.65 44.56
-ASVD 20% 8.07 101.1 11.86 28.8 63.55 62.59 46.12 32.68 73.72 24.65 47.44
-Taco-SVD 20.6% 11.1 56.57 14.76 26.4 60.62 66.22 46.55 35.07 73.11 24.22 47.46
-LLaMA3.1-8B -Instruct Dense 0% 7.21 12.33 12.06 33.8 81.9 73.72 59.1 51.71 79.87 39.46 59.94
-LaCo 19% 753.98 1577.44 807.12 26 48.78 64.56 33.2 29.61 65.07 30.25 42.50
-SliceGPT 20% 52.78 460.73 324.71 14.6 43.48 51.38 29.6 22.78 58.38 21.78 34.57
-ASVD 20% 2443.99 4469.54 2842.23 11.8 28.37 48.62 26.63 19.2 54.84 18.79 29.75
-Taco-SVD 20.1% 37.86 63.97 42.25 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
-Table 1: Zero-shot performance of Taco-SVD and training-free baselines under 20% compression ratio.
-Results are reported on three language modeling datasets (perplexity) and seven reasoning datasets (individual and average accuracy).
-Bold values indicate the best performance.
-Model Method Ratio WikiText-2 PTB C4 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
-Mistral-7B Dense 0% 5.25 23.73 8.36 32.6 80.85 73.88 61.22 50.43 80.58 35.88 59.35
-LoRA 25% 11.56 34.05 14.86 24.6 68.73 70.72 51.18 40.1 71.93 26.37 50.52
-LS-FFN 25.2% 16.12 59.98 18.4 21.8 66.88 68.03 49.92 37.8 70.35 27.2 48.85
-LS-Layer 24.1% 14.67 54.31 16.61 21.6 68.77 70.09 45.02 39.25 72.25 26.83 49.12
-Taco-SVD 24.3% 11.62 44.14 14.85 26.2 69.28 70.32 51.7 40.87 73.07 27.17 51.23
-LLaMA-2-7B Dense 0% 5.47 20.83 7.29 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
-LoRA 25% 11.7 35.39 13.95 26.2 57.79 63.06 48.15 33.28 71.16 24.15 46.26
-LS-FFN 25.7% 13.94 37.43 14.61 24.6 55.98 62.51 48.32 33.28 71 25.19 45.84
-LS-Layer 24% 11.57 32.62 12.91 25.4 59.85 61.88 50.21 35.15 72.91 23.85 47.04
-Taco-SVD 24.2% 9.59 30.38 12.22 28.4 60.52 64.56 50.32 36.01 71.76 24.85 48.06
-LLaMA-2-13B Dense 0% 4.88 28.92 6.99 32.4 73.27 69.61 59.72 45.56 78.78 30.02 55.62
-LoRA 25% 8.46 58.15 11.35 26.6 63.05 67.4 53.47 38.4 73.5 25.33 49.68
-LS-FFN 25.4% 8.11 76.3 11.22 28.4 61.07 67.17 54.04 38.23 73.29 25.49 49.67
-LS-Layer 24.4% 8.23 77.04 11.09 29 62.42 67.25 54.25 39.76 74.21 26.06 50.42
-Taco-SVD 24.2% 7.52 38.31 10.42 28.6 63.3 67.96 54.63 38.74 73.88 27.27 50.63
-LLaMA3.1-8B -Instruct Dense 0% 7.21 12.33 12.06 33.8 81.9 73.72 59.1 51.71 79.87 39.46 59.94
-LoRA 22.3% 14.56 21.24 21.8 27.6 69.65 69.38 50.54 41.3 72.69 30.18 51.62
-LS-FFN 22% 16.79 24.76 24.14 26.4 70.24 70.17 48.97 40.53 72.03 30.25 51.23
-LS-Layer 21.7% 14.27 20.85 20.79 27.2 73.53 69.38 51.22 42.66 73.39 31.56 52.71
-Taco-SVD 21% 14.13 20.99 21.73 28.2 72.01 69.93 51.5 43.26 73.5 31.26 52.81
-Table 2: Performance comparison of Taco-SVD + FT and training-based baselines under 25% compression ratio.
+For all baselines, we randomly sample 512 instances from the WikiText-2 dataset as the representation evaluation dataset.
+For pruning-based with post-training compensation baselines, all models are trained on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32.
+Further experimental details can be found in Appendix A.3.
+Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
+Method
+PPL $\downarrow$
+(WikiText-2)
+PPL $\downarrow$
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+PPL $\downarrow$
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+PPL $\downarrow$
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+PPL $\downarrow$
+Acc $\uparrow$
+Average
+Original 5.25 23.73 0.59 5.47 20.83 0.53 4.88 28.92 0.56 7.21 12.33 0.60
+ShortGPT 23.95 80.70 0.46 18.45 57.58 0.43 27.26 59.14 0.44 140.47 262.47 0.44
+LaCo 244.53 524.26 0.34 39.43 93.98 0.43 38.27 125.98 0.42 753.98 1577.44 0.43
+SliceGPT 9.04 107.65 0.39 8.63 99.25 0.41 7.44 99.77 0.45 52.78 460.73 0.35
+Ours 18.42 77.98 0.46 16.12 44.09 0.44 11.1 56.57 0.47 37.86 63.97 0.47
+Table 1: Perplexity($\downarrow$) of GRASP and pruning-based without post-training compensation baselines on the WikiText-2 and PTB datasets, as well as the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 20% compression ratio.
+The best performance is marked in bold.
+Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
+Method
+PPL $\downarrow$
+(WikiText-2)
+PPL $\downarrow$
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+PPL $\downarrow$
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+PPL $\downarrow$
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+PPL $\downarrow$
+Acc $\uparrow$
+Average
+Original 5.25 23.73 0.59 5.47 20.83 0.53 4.88 28.92 0.56 7.21 12.33 0.60
+LoRA 11.56 34.05 0.51 11.7 35.39 0.46 8.46 58.15 0.50 14.56 21.24 0.52
+LS-FFN 16.12 59.98 0.49 13.94 37.43 0.46 8.11 76.3 0.50 16.79 24.76 0.51
+LS-Layer 14.67 54.31 0.49 11.57 32.62 0.47 8.23 77.04 0.50 14.27 20.85 0.53
+Ours* 11.62 44.14 0.51 9.59 30.38 0.48 7.52 38.31 0.51 14.13 20.99 0.53
+Table 2: Perplexity($\downarrow$) of GRASP (Ours* denotes the version with post-training compensation) and pruning-based with post-training compensation baselines on the WikiText-2 and PTB datasets, as well as the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 25% compression ratio.
 LS-FFN and LS-Layer denote two variants of LLM-Streamline.
-Results are reported on three language modeling datasets (perplexity) and seven reasoning datasets (individual and average accuracy).
-Bold values indicate the best performance.
-4.2 Results
-We conducted comprehensive experiments to evaluate Taco-SVD from four key perspectives: (a) zero-shot compression performance on language generation and reasoning tasks, (b) effectiveness as an initialization method compared to training-based approaches, (c) robustness at high compression ratios, and (d) compression efficiency and inference speed on real hardware.
-Training-Free Compression
-To evaluate the generalizability of Taco-SVD, we compared its performance with training-free baselines under approximately 20% compression ratio across four models with varying scales and architectures: Mistral-7B, LLaMA2-7B, LLaMA2-13B, and LLaMA3.1-8B-Instruct.
-The results, summarized in Table 1, demonstrate that Taco-SVD consistently achieves strong performance across all models and tasks, maintaining competitive perplexity and accuracy even on challenging architectures like LLaMA3.
-Notably, Taco-SVD exhibits significantly improved robustness compared to baselines, effectively addressing the variability in sensitivity to pruning across diverse architectures.
-Training-based Compression
-In this section, we further demonstrate that Taco-SVD can provide better initialization for fine-tuning, enabling faster and more accurate recovery of the compressed model compared to other methods.
-To ensure that the number of trainable parameters remains approximately the same across all methods, we retain only 10% of the parameters in each redundant layer and set only the parameters within these redundant layers to be trainable.
-As shown in Table 2, Taco-SVD consistently outperforms LoRA and LLM-Streamline across different LLMs.
-In addition, we plot the validation loss curves during the training process for each method in the Appendix A.4, we can observe that Taco-SVD has better initialization, leading to more stable and faster convergence during fine-tuning.
-Performance at High Compression Ratio
-To evaluate the performance and stability of our proposed method at high compression ratios, we compare Taco-SVD with the baselines on LLaMA2-7B at compression ratios of 25%, 30%, 35% and 40%.
-The perplexity (PPL) on WikiText-2, PTB and the average accuracy across all reasoning benchmarks are shown in Figure 3, with detailed results for each benchmark presented in the Appendix A.5.
-We make the following three observations: (a) Most layer removal methods cannot tolerate high compression ratios, with performance dropping sharply at higher compression levels.
-(b) Compared to training-free baselines, Taco-SVD consistently achieves better accuracy across all compression ratios.
-(c) Fine-tuning Taco-SVD significantly recovers the accuracy of compressed models, even at high compression ratios.
-Figure 3: Zero-shot performance of LLaMA2-7B compressed by Taco-SVD and baselines under 20% to 40% compression ratio.
-Top: Perplexity on WikiText-2 (left) and PTB (right).
-Bottom: Average accuracy on seven reasoning benchmarks.
-Compression Costs and Acceleration Effect
-Taco-SVD can not only compress LLMs at very low cost but also improve inference efficiency on real hardware.
-To demonstrate this, we measure the compression time and peak memory consumption of Taco-SVD when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio.
-These metrics, summarized in Table 3, indicate that Taco-SVD achieves a good balance between compression time and memory consumption.
-Additionally, to assess the acceleration effect of Taco-SVD, we compared the throughput (tokens generated per second) of original LLaMA2-7B and its compressed version by layer removal and Taco-SVD with different sequence lengths and batch sizes.
-As shown in Figure 4, Taco-SVD consistently ensures an enhancement in the generation speed and achieves an acceleration effect comparable to pure layer removal.
-It is important to note that although Taco-SVD retains a small portion of parameters in the redundant layers to mitigate the internal inconsistencies introduced by layer removal, this structure is extremely low-rank, resulting in negligible inference latency while maintaining task performance.
+The best performance is marked in bold.
+Ratio Method WikiText-2 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA GSM8K Average Percentage
+0% Original 5.68 0.28 0.67 0.67 0.56 0.38 0.78 0.27 0.09 0.46 100.0
+20% FWSVD 1727 0.15 0.31 0.50 0.26 0.23 0.56 0.21 0.00 0.28 60.8
+ASVD 11.14 0.25 0.53 0.64 0.41 0.27 0.68 0.24 0.04 0.38 82.6
+SVD-LLM 7.94 0.22 0.58 0.63 0.43 0.29 0.69 0.24 0.05 0.39 84.7
+Ours 14.79 0.22 0.52 0.64 0.43 0.32 0.70 0.24 0.03 0.39 84.7
+Ours* 10.19 0.24 0.59 0.63 0.5 0.35 0.73 0.25 0.04 0.42 91.3
+30% FWSVD 20127 0.17 0.26 0.49 0.26 0.22 0.51 0.19 0.00 0.26 56.5
+ASVD 51 0.18 0.43 0.53 0.37 0.25 0.65 0.21 0.00 0.33 71.7
+SVD-LLM 9.56 0.20 0.48 0.59 0.37 0.26 0.65 0.22 0.03 0.35 76.1
+Ours 30.48 0.19 0.42 0.62 0.39 0.28 0.64 0.23 0.02 0.35 76.1
+Ours* 12.63 0.24 0.54 0.64 0.46 0.32 0.69 0.24 0.04 0.40 87.0
+40% FWSVD 18156 0.16 0.26 0.51 0.26 0.22 0.53 0.21 0.00 0.27 58.7
+ASVD 15358 0.13 0.28 0.48 0.26 0.22 0.55 0.19 0.00 0.26 56.5
+SVD-LLM 13.11 0.19 0.42 0.58 0.33 0.25 0.60 0.21 0.02 0.33 71.7
+Ours 62.34 0.18 0.37 0.57 0.35 0.27 0.61 0.21 0.01 0.32 69.6
+Ours* 16.25 0.22 0.49 0.63 0.43 0.3 0.68 0.23 0.02 0.38 82.6
+50% FWSVD 24391 0.12 0.26 0.50 0.26 0.23 0.53 0.20 0.00 0.26 56.5
+ASVD 15358 0.12 0.26 0.51 0.26 0.22 0.52 0.19 0.00 0.26 56.5
+SVD-LLM 23.97 0.16 0.33 0.54 0.29 0.23 0.56 0.21 0.00 0.29 63.0
+Ours 102.58 0.13 0.29 0.53 0.28 0.23 0.53 0.20 0.01 0.28 60.9
+Ours* 23.69 0.18 0.4 0.56 0.35 0.26 0.61 0.21 0.02 0.32 69.6
+[…]
+Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
+Method
+PPL $\downarrow$
+(WikiText-2)
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+Acc $\uparrow$
+Average
+Original 5.25 0.59 5.68 0.52 5.47 0.53 7.21 0.60
+FWSVD 6357 0.32 1727 0.32 2360 0.31 3256.7 0.29
+ASVD 19.28 0.4 11.14 0.44 9.70 0.46 2443.99 0.30
+SVD-LLM 10.21 0.41 7.94 0.44 8.50 0.47 - -
+Ours 18.42 0.45 14.79 0.44 16.12 0.44 37.86 0.47
+Ours* 11.62 0.51 10.19 0.47 9.59 0.48 14.13 0.53
+Table 4: Perplexity($\downarrow$) of GRASP and SVD-based baselines on the WikiText-2 datasets and the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 20% compression ratio.
+"†" indicates that we refer to the results in the original paper.
+The best performance is marked in bold.
 Model
+Single-Doc
+QA
+Multi-Doc
+QA
+Summarization
+Task
+Few-shot
+Learning
+Synthetic
+Task
+Code
+Completion
 Average
-Compression Time(h)
-Memory
-Consumption(MiB)
-LaCO 0.05 45046
-ASVD 5.5 17632
-SliceGPT 0.6 18195
-Taco-SVD 0.16 29050
-Table 3: Compression time and peak memory usage of Taco-SVD and baselines for LLaMA2-7B under 25% compression ratio on a single A100 GPU.
-Figure 4: Throughput of LLaMA2-7B and Taco-SVD compressed model under 25% compression ratio on a single A100 GPU.
+LLaMA3.1-8B-Instruct 12.26 16.16 22.19 42.52 9.49 53.91 26.09
+GRASP* 21.69 24.41 20.03 45.84 9.67 35.92 26.26
+Table 5: Evaluations of LLaMA3.1-8B-Instruct and its compressed version of GRASP* under 20% compression ratio on LongBench benchmarks.
+The average results across six task categories are shown in this table, with additional details provided in the Appendix A.7.
+4.2 Comparison with Pruning-based LLM Compression Methods
+Post-training compensation plays a crucial role in restoring the performance of compressed LLMs.
+To ensure a fair comparison, we compare GRASP with two categories of pruning-based compression methods.
+(1) Pruning without post-training compensation, including two layer removal methods, ShortGPT (Men et al., 2024), LaCo (Yang et al., 2024), and state-of-the-art structured pruning method: SliceGPT (Ashkboos et al., 2024).
+(2) Pruning with post-training compensation: ShortGPT with LoRA (Hu et al., 2022) and LLM-Streamline (Chen et al., 2024).
+We provide a detailed comparison of these pruning-based LLM compression methods in Appendix A.4.
+Pruning without post-training compensation
+We present the perplexity on WikiText-2 and PTB datasets, as well as the average accuracy on common sense reasoning benchmarks for different methods in Table 1.
+The results demonstrate that GRASP consistently achieves strong performance across all models and tasks.
+Specifically, in common sense reasoning tasks, GRASP surpasses SliceGPT by 15% in accuracy on Mistral-7B and by 34% on LLaMA3.1-8B-Instruct.
+For language generation tasks, while SliceGPT achieves lower perplexity on the WikiText-2 dataset, GRASP consistently outperforms the baselines on the PTB dataset.
+We hypothesize this discrepancy arises from the potential overfitting problem of baselines to the calibration dataset.
+Notably, GRASP exhibits significantly improved robustness compared to baselines, effectively addressing the variability in sensitivity to pruning across diverse models.
+Detailed results for each common sense reasoning benchmark are provided in Appendix A.5.
+Pruning with Post-Training Compensation
+We further demonstrate that GRASP provides a better initialization for post-training compensation, enabling more accurate recovery of the compressed model compared to other methods.
+To ensure a fair comparison, we constrain the number of trainable parameters to remain approximately the same across all methods by retaining only 10% of the parameters in each redundant layer and allowing only these parameters to be trainable.
+As shown in Table 2, GRASP consistently outperforms LLM-Streamline across different LLMs.
+Furthermore, we present the validation loss curves for each method during post-training in Appendix A.6, which show that GRASP GRASP enables more stable and faster convergence.
+More Detailed results are provided in Appendix A.5.
+4.3 Comparison with SVD-based LLM Compression Methods
+In this section, we compare the performance of GRASP with state-of-the-art SVD-based LLM compression methods—FWSVD (Hsu et al., 2022), ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024)—from two distinct perspectives: (1) performance under different compression ratios and (2) performance across different LLMs.
+Performance under Different Compression Ratios.
+To evaluate the performance and stability of our proposed method under various compression ratios, we compare GRASP with the SVD-based baselines on LLaMA-7B under compression ratios ranging from 20% to 50%.
+Table 3 summarizes the results for different methods.
+The results demonstrate that our proposed GRASP consistently outperforms the baseline methods on most benchmarks, except for a slightly higher perplexity on WikiText-2 compared to SVD-LLM.
+Specifically, GRASP retains more than 91% of the original performance at a 20% compression ratio and 87% under the compression ratio of 30%.
+More importantly, with fast and resource-efficient post-training, GRASP enables rapid accuracy recovery, achieving 70% of the original model performance even at a 50% compression ratio.
+Performance on different LLMs.
+To evaluate the generalizability of GRASP across different LLMs, we compare its performance with the SVD-based baselines under 20% compression ratio on four models from two distinct LLM families-LLaMA 2-7B, LLaMA 2-13B, LLaMA 3.1-8B-Instruct and Mistral-7B.
+The results, presented in Table 4, indicate that while existing methods achieve competitive performance on the LLaMA 2 series, they exhibit a sharp performance drop on other architectures, particularly on LLaMA 3.
+In contrast, GRASP consistently outperforms baseline methods across all architectures and demonstrates greater stability across different LLMs.
+4.4 Performance in Long-Form Text Processing
+In this section, we evaluate the performance of GRASP on LongBench (Bai et al., 2023) to assess its performance on long-form reasoning and complex generative tasks.
+The average performance across six task categories: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion are summarized in Table 5, with detailed results provided in the Appendix A.7.
+As shown, the compressed version by GRASP under 20% compression ratio still maintains strong performance and even surpasses the original LLaMA3.1-8B-Instruct model in single-document QA, multi-document QA, and few-shot learning tasks.
+This demonstrates the effectiveness of GRASP in rapidly recovering accuracy and generative capability through post-training compensation after pruning.
+4.5 Compression Costs and Inference Efficiency of GRASP
+GRASP can compress LLMs at very low cost while improve the inference efficiency on real hardware.
+To evaluate its acceleration effect, we compared the throughput (tokens generated per second) of original LLaMA2-7B and its compressed version by GRASP with different sequence lengths and batch sizes.
+As shown in Figure 3, GRASP consistently consistently improves generation speed and achieves an acceleration effect comparable to pure layer removal.
+It is important to note that although GRASP retains a small portion of parameters in the redundant layers to mitigate the internal inconsistencies introduced by layer removal, this structure is extremely low-rank, resulting in negligible inference latency while maintaining task performance.
+Additionally, we also measure the compression time and peak memory consumption of GRASP when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio.
+The detailed results are provided in Appendix A.8.
+Figure 3: Throughput of LLaMA2-7B and GRASP compressed model under 25% compression ratio on a single A100 GPU.
 Top: Throughput across different sequence lengths (batch size = 32).
@@
 Calibration Data
-We conducted ablation studies to investigate the impact of calibration dataset and the number of calibration data used for singular value gradient attribution.
-Table 4 summarizes the results of Taco-SVD when compressing LLaMA3.1-8B-Instruct under 20% compression ratio using different calibration datasets and varying numbers of calibration data from WikiText-2.
-The results demonstrate that Taco-SVD is robust to both the choice of calibration dataset and the quantity of calibration data.
-Notably, even with as few as 64 calibration samples, Taco-SVD achieves competitive performance, highlighting its efficiency in leveraging minimal data for effective compression.
-We report the detailed results in the Appendix A.6.
+We conducted ablation studies to investigate the impact of representation evaluation dataset and the number of data used for singular value gradient attribution.
+Table 6 summarizes GRASP’s performance when compressing LLaMA3.1-8B-Instruct using different representation evaluation datasets and varying numbers of data from WikiText-2.
+The results demonstrate that GRASP is robust to the choice of both the representation evaluation dataset and the data number.
+Notably, GRASP achieves competitive performance with just 64 samples, highlighting its ability to perform effective compression with minimal data.
+Detailed results can be found in Appendix A.9.
 Ablation Type
@@
 512 37.86 63.97 47.12
-Table 4: Comparison of Taco-SVD using different types and amounts of calibration data for compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
+Table 6: Comparison of GRASP using different types and amounts of data for compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
 One-shot Pruning vs Iterative Pruning
-As detailed in Section 3.4, Taco-SVD employs a layer-wise backward compression strategy to compress redundant layers.
+As detailed in Section 3.4, GRASP employs a layer-wise backward compression strategy to compress redundant layers.
 This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers.
-Table 5 shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
-Pruning Strategy WikiText-2 PTB
-Average
-Accuracy
-Compression
-Time(h)
-One-shot Pruning 37.86 63.97 47.12 0.16
-Iterative Pruning 38.39 72.18 47.13 0.22
-Table 5: Comparison of one-shot pruning and iterative pruning for LLaMA3.1-8B-Instruct under 20% compression ratio.
+We present the ablation results in Table 12 in Appendix A.10, which shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
 5 Discussion
-5.1 How Taco-SVD Stabilizes Layer Removal and Enhances Model Consistency
-Instabilities Caused by Layer Removal
-Layer removal can disrupt the internal consistency of a model, leading to performance degradation.
-In response, the model attempts to compensate for the loss of removed layers, triggering a process we refer to as self-repair (Rushing and Nanda, 2024).
-While this self-repair mechanism helps the model maintain partial functionality, it is often noisy and incomplete.
-The process cannot fully recover the removed components, resulting in higher perplexity and reduced cross-task generalization.
-This phenomenon is task-dependent, with certain tasks being more sensitive to the loss of critical task-relevant transformations.
-Taco-SVD as a Stabilizing Solution
-Taco-SVD effectively addresses these issues by preserving key task-critical singular value directions, reducing the need for self-repair and stabilizing the model’s internal consistency.
-Specifically, Taco-SVD:
-  • Retains task-critical singular values, minimizing disruption to the residual stream.
-This helps maintain the original structure and function of the model, reducing perplexity and improving performance across diverse tasks.
-  • Mitigates the reliance on self-repair, ensuring that the model does not need to rely on incomplete compensations.
-By aligning the retained singular values with task-specific objectives, Taco-SVD facilitates smoother adjustments that preserve both task performance and internal consistency.
-5.2 Taco-SVD vs LoRA vs LLM-Streamline
-Limitations of fine-tuning Pruned LLMs with LoRA
-Parameter-efficient fine-tuning methods such as LoRA introduce additional parameters across the remaining layers to compensate for the removed layers.
-However, training the pruned model with LoRA is a process of redistributing the function of the pruned layers (Chen et al., 2024), these updates often introduce intruder dimensions, which may lead to potential issues such as Spectral Dilution (Shuttleworth et al., 2024): LoRA’s weight updates alter the original singular value spectrum, the new singular vectors added by LoRA are often orthogonal to the original pre-trained singular vectors, which diminishes the contribution of task-agnostic features learned during pre-training.
-Limitations of Layer Replacement with Dense Modules
-Methods like LLM-Streamline (Chen et al., 2024) train a dense module from scratch to approximate the functionality of redundant layers within LLMs.
-However, dense modules possess high-dimensional internal structures, which may cause overfitting in scenarios with limited data.
-Taco-SVD’s Generalization Advantage via Low-Rank Learning
-In contrast, Taco-SVD avoids such issues by preserving the original spectral structure while aligning it with task-specific objectives.
-  • Taco-SVD identifies and retains the singular value directions in the redundant layer that are most critical to task performance.
-It maintains the original pre-trained singular value spectrum, preventing the introduction of orthogonal updates to the original singular vectors in the remaining layers.
-  • Previous literature has demonstrated the effectiveness and practicality of low-rank learning in deep networks (Oymak et al., 2019; Kamalakara et al., 2022; Yaras et al., 2024).
-Therefore, instead of training a dense module from scratch, Taco-SVD fine-tunes only the minimal task-aware subspaces in redundant layers.
-It can avoid overfitting by constraining the model to learn within a low-rank structure.
+GRASP preserves representation stability.
+Layer removal disrupts the internal spectral structure of LLMs, often introducing inconsistencies in the learned representations.
+These disruptions lead to degradation in perplexity and loss of linguistic coherence, as the model struggles to compensate for missing transformations.
+GRASP stabilizes the compression process by selectively retaining representation-critical singular values, ensuring that the model’s residual stream remains structurally consistent.
+Unlike direct layer removal, which forces the model to undergo self-repair (Rushing and Nanda, 2024), GRASP proactively mitigates this instability by preserving the most influential spectral components, thereby reducing the burden on later layers to reconstruct lost information.
+GRASP accelerates adaptation via low-rank learning.
+Fine-tuning pruned models using parameter-efficient methods such as LoRA often introduces intruder dimensions into the pre-trained singular spectrum (Shuttleworth et al., 2024), leading to spectral dilution that weakens the contribution of pre-trained features.
+Similarly, replacing layers with dense modules, as in LLM-Streamline (Chen et al., 2024), requires training high-dimensional structures from scratch, which increases the risk of overfitting when data is limited.
+In contrast, GRASP identifies and fine-tunes the most essential low-rank subspaces within redundant layers, preserving the original spectral alignment while adapting efficiently to new tasks.
+Prior research has established the effectiveness of low-rank learning in deep networks (Oymak et al., 2019; Kamalakara et al., 2022), and GRASP leverages this principle to optimize task-specific performance while maintaining pre-trained feature integrity.
 6 Conclusion
-We propose Taco-SVD, a task-aware framework for compressing large language models by retaining task-critical singular value directions.
-Taco-SVD mitigates the disruptions caused by traditional layer removal and avoids the need for computationally expensive fine-tuning compensation by preserving key task-critical transformations.
-Experimental results show that Taco-SVD achieves lower perplexity, better task performance, and improved cross-task robustness with minimal resource overhead.
-This work highlights the effectiveness of aligning low-rank decompositions with task-specific objectives for efficient and robust model compression.
+We present GRASP, a novel framework for compressing large language models by selectively retaining representation-critical singular value directions.
+By mitigating the disruptions caused by layer removal, GRASP preserves essential transformations necessary for maintaining model stability and generalization.
+Experimental results demonstrate that GRASP achieves efficient and robust compression, improving perplexity and downstream task performance compared to existing methods while maintaining minimal computational overhead.
 7 Limitations
-While Taco-SVD demonstrates strong performance across various tasks and architectures, we acknowledge certain limitations that highlight opportunities for future research:
-Dependence on Calibration Data Taco-SVD relies on calibration data for singular value gradient attribution.
-In our experiments, using as few as 64 calibration samples from the general-purpose WikiText dataset achieved strong performance across multiple tasks.
-However, while this approach works well in these settings, we have not yet explored whether using domain-specific calibration data could yield better results in specialized domains.
+While GRASP demonstrates strong performance across various tasks and architectures, we acknowledge certain limitations that highlight opportunities for future research:
+Dependence on Calibration Data GRASP relies on representation evaluation dataset for singular value gradient attribution.
+In our experiments, using as few as 64 samples from the WikiText dataset achieved strong performance across multiple tasks.
+However, while this approach works well in these settings, we have not yet explored whether using domain-specific dataset could yield better results in specialized domains.
 Further testing with such data may provide valuable insights into its potential advantages.
-Trade-Off Between Compression and Retained Parameters To ensure robustness and task consistency, Taco-SVD retains a small fraction of low-rank parameters in redundant layers, mitigating disruptions caused by pure layer removal.
+Trade-Off Between Compression and Retained Parameters To ensure robustness and task consistency, GRASP retains a small fraction of low-rank parameters in redundant layers, mitigating disruptions caused by pure layer removal.
 This design choice introduces a slight trade-off in terms of parameter retention, which could limit its suitability for applications requiring extreme compression ratios.
-Further reduction techniques may complement Taco-SVD in such scenarios.
-Scalability to Extremely Large Models While Taco-SVD performs effectively on models like LLaMA2 and LLaMA3, its application to ultra-large-scale models (e.g., hundreds of billions of parameters) could pose additional computational challenges.
+Further reduction techniques may complement GRASP in such scenarios.
+Scalability to Extremely Large Models While GRASP performs effectively on models like LLaMA2 and LLaMA3, its application to ultra-large-scale models (e.g., hundreds of billions of parameters) could pose additional computational challenges.
 Future work could focus on optimizing the algorithm for such architectures, enabling its application to next-generation LLMs.
@@
 All datasets used in this work are publicly available and have been appropriately cited, ensuring compliance with data usage agreements and privacy regulations.
-While Taco-SVD is designed to optimize the efficiency and scalability of large language models, we recognize that such technologies could be misused in applications that may perpetuate harmful biases or deploy models in contexts lacking adequate oversight.
+While GRASP is designed to optimize the efficiency and scalability of large language models, we recognize that such technologies could be misused in applications that may perpetuate harmful biases or deploy models in contexts lacking adequate oversight.
 To mitigate these risks, we advocate for responsible deployment practices, including thorough testing and monitoring for unintended biases.
@@
 $\frac{\partial\sigma_{i}}{\partial W}=u_{i}v_{i}^{T}$
-For a calibration dataset $D$, the gradient of a singular value $\sigma_{i}$ with respect to the task loss can be interpreted as the projection of the weight gradient matrix $G$ onto the corresponding singular direction, given by:
+For a representation evaluation dataset $D$, the gradient of a singular value $\sigma_{i}$ with respect to the task loss can be interpreted as the projection of the weight gradient matrix $G$ onto the corresponding singular direction, given by:
 $\frac{\partial L}{\partial\sigma_{i}}=u_{i}^{T}\frac{\partial L}{\partial W}v_{i}$
@@
 This section provides the implementation of Layer-wise Low-Rank Decomposition, as detailed in Algorithm 1.
-Algorithm 1 Task-Aware SVD for LLM Compression
-1:Model $M$ with $N$ layers, Calibration Dataset $D$, Retain Ratio $r$
+Algorithm 1 Gradient-Guided SVD for LLM Compression
+1:Model $M$ with $N$ layers, Representation Evaluation Dataset $D$, Retain Ratio $r$
 2:$\text{W}\in\{W_{q},W_{k},W_{v},W_{o},W_{up},W_{down}\}$ for weight matrices in self-attention and MLP blocks
@@
 8:Select $L$ layers with highest cosine similarity.
-9:Step 2: Task-aware SVD Compression:
+9:Step 2: Gradient-guided SVD Compression:
 10:for each layer $l\in L$, starting from the last layer do
@@
 17:return $\tilde{M}$
-Figure 5: Validation loss curves of different methods on the Alpaca dataset (1 epoch)
+Method Metric Calibration Data Need Post-Training Training Data Training Dataset Size Training Module
+ShortGPT Cosine Similarity WikiText-2 No None None None
+LaCo Cosine Similarity WikiText-2 Optional Unpublished 1B Full Parameters
+SliceGPT PCA
+WikiText-2
+Alpaca
+Optional Alpaca 5k Full Parameters
+Shortened LLaMA
+Taylor
+Perplexity
+BookCorpus Optional
+SlimPajama
+Alpaca
+627B
+50k
+Full Parameters
+LoRA-Adapter
+LLM-Streamline Cosine Similarity WikiText-2 Yes SlimPajama 30k
+Lightweight
+Network
+GRASP Cosine Similarity WikiText-2 Optional Alpaca 50k
+Low-rank
+Modules
+Table 7: Comparison of pruning-based LLM compression methods, where the metric indicates the criterion used to identify redundant modules.
+"Optional" refers to methods that can either work without post-training or recover performance through post-training.
+Shortened LLaMA consists of two training stages: initial continual pre-training on the SlimPajama dataset, followed by LoRA fine-tuning on the Alpaca dataset.
 A.3 Experimental Setup and Hyperparameters Configuration
-To ensure a fair comparison, all experimental setups are consistent across all methods. Below we describe the experimental setup and hyperparameters configuration in detail.
-Training-free Setting
-We report the results of Taco-SVD and other training-free baselines under the compression ratio of 20% in Table 1.
-Specifically, we selected the same 256 samples from the WikiText-2 dataset as calibration data.
-Training-based Setting
-Table 6 provides the detailed configurations of training-based methods under the compression ratio of 20%. All our experiments are conducted on a single NVIDIA A100 GPU with mixed precision enabled.
+To ensure a fair comparison, all experimental setups are consistent across all methods. In the following, we describe the experimental setup and hyperparameters configuration in detail.
+Metric
+For the WikiText-2 and PTB datasets, we use perplexity ($\downarrow$) to assess model’s generative capacity.
+For the seven common sense reasoning datasets, performance is measured by accuracy ($\uparrow$), and for the GSM8K dataset, performance is evaluated using Exact Match Accuracy ($\uparrow$).
+All common sense reasoning tasks are tested in zero-shot setting using the LM-Evaluation-Harness framework (Gao et al., 2024), while GSM8K is tested in the five-shot setting.
+Hyperparameters Configurations
+For post-training compensation, all models compressed by different methods are trained on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32.
+We use AdamW (Loshchilov and Hutter, 2019) as our optimizer and set the learning rate to $3\times 10^{-4}$.
+All our experiments are conducted on a single A100 GPU with mixed precision enabled.
+Table 8 provides the detailed configurations of post-training compensation.
 HyperParameters Setting
@@
 Prompt-Template Alpaca Template
-Table 6: Experimental setup and hyperparameter configurations.
-Method Ratio WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
-Dense 0% 5.47 20.83 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
-LaCo 18% 39.43 93.98 26.2 48.36 59.19 41.93 32.42 68.72 24.02 42.98
-SliceGPT 21.50% 8.63 99.25 22 54.34 60.62 37.4 27.99 63.38 23.18 41.27
-ASVD 20% 9.7 88.39 27.8 58.41 61.09 46.97 32.42 71.05 25.09 46.12
-Taco-SVD 21.60% 16.12 44.09 24 53.59 62.9 42.94 32.76 71.06 22.88 44.30
-LaCo 24% 50.33 102.84 25 45.33 60.06 38.3 29.69 65.51 23.99 41.13
-SliceGPT 25.10% 9.54 118.17 20.6 50.91 59.63 35.66 26.41 61.62 23.43 39.75
-ASVD 25% 53.48 374.45 17 39.35 53.12 32.64 22.78 61.48 22.95 35.62
-Taco-SVD 25% 30.2 60.74 21.6 46.59 61.8 40.41 28.84 66.1 23.18 41.22
-LaCo 30% 94.5 122.09 22 42.26 59.43 35.7 28.41 62.89 24.32 39.29
-SliceGPT 29.10% 10.91 145.24 18.8 47.56 58.64 34.18 25.34 59.58 23.15 38.18
-ASVD 30% 1877.63 3340.65 14.4 27.78 50.59 26.37 21.84 51.58 20.88 30.49
-Taco-SVD 30% 35.38 88.96 19.4 42.72 61.17 38.43 27.73 63.98 23.18 39.52
-LaCo 36% 138.82 164.02 19 38.34 58.96 33.71 27.13 60.55 23.55 37.32
-SliceGPT 33.20% 12.8 176.54 17.8 44.23 57.61 32.28 23.55 58.59 22.68 36.68
-ASVD 35% 2599.79 6529.72 13.6 26.81 50.83 25.94 23.12 53.59 19.33 30.46
-Taco-SVD 35.00% 68.43 125.79 18.8 40.25 60.35 36.77 27.18 61.11 22.63 38.16
-LaCo 39% 386.55 289.98 18.4 33.96 56.99 31.34 26.88 58.6 23.45 35.66
-SliceGPT 41.40% 19.1 257.66 15.8 35.94 53.43 29.62 20.82 55.98 22.01 33.37
-ASVD 40% NaN NaN 13.4 25.76 51.3 26.1 21.67 54.19 20.64 30.44
-Taco-SVD 40.10% 80.42 247.55 18 36.83 59.59 33.31 26.62 59.63 21.88 36.55
-Table 7: Zero-shot performance of LLaMA2-7B compressed by Taco-SVD and training-free baselines under 20% to 40% compression. Bold values indicate the best performance.
-A.4 Validation Loss Curves
+Table 8: Experimental setup and hyperparameters configurations.
+A.4 Comparison of Concurrent Pruning-based Methods
+We provide a detailed comparison of concurrent pruning-based LLM compression methods, and the results are summarized in Table 7.
+A.5 Detailed Results on Common Sense Reasoning Benchmarks
+In this section, we provide the detailed results for each benchmark.
+Table 9 summarizes the results of pruning-based methods without post-training compensation on the common sense reasoning benchmark, while Table 10 presents the results of methods with post-training compensation.
+Figure 4: Validation loss curves of different methods on the Alpaca dataset (1 epoch)
+Model Method Ratio Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
+Mistral-7B-v0.1 Dense 0.0% 0.33 0.81 0.74 0.61 0.50 0.81 0.36 0.59
+LaCo 21.1% 0.20 0.35 0.58 0.26 0.25 0.53 0.24 0.34
+ShortGPT 21.1% 0.19 0.57 0.68 0.46 0.37 0.71 0.26 0.46
+SliceGPT 20.0% 0.19 0.51 0.59 0.35 0.25 0.61 0.23 0.39
+GRASP 20.0% 0.21 0.56 0.68 0.43 0.38 0.67 0.26 0.46
+LLaMA2-7B Dense 0.0% 0.32 0.69 0.67 0.57 0.40 0.78 0.28 0.53
+LaCo 18.1% 0.26 0.48 0.59 0.42 0.32 0.69 0.24 0.43
+ShortGPT 21.1% 0.23 0.49 0.63 0.42 0.31 0.68 0.23 0.43
+SliceGPT 21.5% 0.22 0.54 0.61 0.37 0.28 0.63 0.23 0.41
+GRASP 21.6% 0.24 0.54 0.63 0.43 0.33 0.71 0.23 0.44
+LLaMA-2-13B Dense 0.0% 0.32 0.73 0.70 0.60 0.46 0.79 0.30 0.56
+LaCo 19.5% 0.28 0.52 0.63 0.43 0.33 0.70 0.25 0.45
+ShortGPT 22.1% 0.24 0.50 0.63 0.46 0.33 0.7 0.24 0.44
+SliceGPT 20.0% 0.29 0.59 0.65 0.39 0.32 0.64 0.24 0.45
+GRASP 20.0% 0.26 0.61 0.66 0.47 0.35 0.73 0.24 0.47
+LLaMA3.1-8B-Instruct Dense 0.0% 0.34 0.82 0.74 0.59 0.52 0.80 0.39 0.60
+LaCo 19.0% 0.26 0.49 0.65 0.33 0.30 0.65 0.30 0.42
+ShortGPT 21.7% 0.21 0.57 0.66 0.42 0.32 0.67 0.26 0.44
+SliceGPT 20.0% 0.15 0.43 0.51 0.30 0.23 0.58 0.22 0.35
+GRASP 20.0% 0.22 0.60 0.70 0.44 0.37 0.69 0.28 0.47
+Table 9: Zero-shot performance of GRASP and pruning-based without post-training baselines under 20% compression ratio.
+Results are reported on seven reasoning datasets (individual and average accuracy).
+Bold values indicate the best performance.
+Model Method Ratio Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
+Mistral-7B Dense 0.0% 0.33 0.81 0.74 0.61 0.50 0.81 0.36 0.59
+LoRA 25.1% 0.25 0.69 0.71 0.51 0.40 0.72 0.26 0.51
+LS-FFN 25.2% 0.22 0.67 0.68 0.50 0.38 0.70 0.27 0.49
+LS-Layer 24.1% 0.22 0.69 0.70 0.45 0.39 0.72 0.27 0.49
+GRASP 24.3% 0.26 0.69 0.70 0.52 0.41 0.73 0.27 0.51
+LLaMA-2-7B Dense 0.0% 0.32 0.69 0.67 0.57 0.40 0.78 0.28 0.53
+LoRA 25.0% 0.26 0.58 0.63 0.48 0.33 0.71 0.24 0.46
+LS-FFN 25.7% 0.25 0.56 0.63 0.48 0.33 0.71 0.25 0.46
+LS-Layer 24.0% 0.25 0.60 0.62 0.50 0.35 0.73 0.24 0.47
+GRASP 24.3% 0.28 0.61 0.65 0.50 0.36 0.72 0.25 0.48
+LLaMA-2-13B Dense 0.0% 0.32 0.73 0.70 0.60 0.46 0.79 0.30 0.56
+LoRA 25.0% 0.27 0.63 0.67 0.53 0.38 0.74 0.25 0.50
+LS-FFN 25.4% 0.28 0.61 0.67 0.54 0.38 0.73 0.25 0.50
+LS-Layer 24.4% 0.29 0.62 0.67 0.54 0.40 0.74 0.26 0.50
+GRASP 24.2% 0.29 0.63 0.68 0.55 0.39 0.74 0.27 0.51
+LLaMA3.1-8B-Instruct Dense 0.0% 0.34 0.82 0.74 0.59 0.52 0.80 0.39 0.60
+LoRA 22.3% 0.28 0.70 0.69 0.51 0.41 0.73 0.30 0.52
+LS-FFN 22.0% 0.26 0.70 0.70 0.49 0.41 0.72 0.30 0.51
+LS-Layer 21.7% 0.27 0.74 0.69 0.51 0.43 0.73 0.32 0.53
+GRASP 21.0% 0.28 0.72 0.70 0.52 0.43 0.74 0.31 0.53
+Table 10: Performance comparison of GRASP and pruning-based with post-training baselines under 25% compression ratio.
+LS-FFN and LS-Layer denote two variants of LLM-Streamline.
+Results are reported on seven reasoning datasets (individual and average accuracy).
+Bold values indicate the best performance.
+A.6 Validation Loss Curves
 We plot the validation loss curves when training pruned Mistral-7B with 20% compression ratio with different compensation methods.
-As shown in Figure 5, we can observe that Taco-SVD has better initialization, leading to more stable and faster convergence during fine-tuning.
-A.5 Results under High Compression Ratio
-In this section, we present the results of Taco-SVD compared with the training-free baselines using LLaMA2-7B under high compression ratios, as shown in Table 7.
-The results demonstrate that our proposed method consistently outperforms the baselines, even at high compression ratios of up to 40%.
-A.6 Robustness of Taco-SVD towards different Calibration Dataset
-In this section, we provide details of the ablation studies conducted to investigate the impact of calibration datasets and the number of calibration data used for singular value gradient attribution.
-Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of Taco-SVD when compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
-Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of Taco-SVD to the change in the number of calibration data.
+As shown in Figure 4, we can observe that GRASP has better initialization, leading to more stable and faster convergence during training.
+A.7 Evaluation Results on LongBench
+In this section, we present the detailed results of LLaMA3.1-8B-Instruct and its compressed version by GRASP under 20% compression ratio on LongBench, which are presented in Table 13.
+The results illustrate that GRASP with post-training compensation still maintains superior performance on long-form reasoning and complex generative tasks.
+A.8 Compression Costs of GRASP
+To assess the compressed costs of GRASP, we measure its compression time and peak memory consumption when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio.
+These metrics, summarized in Table 11, indicate that GRASP achieves a good balance between compression time and memory consumption.
+Model
+Average
+Compression Time(h)
+Compression Memory
+Consumption(MiB)
+LaCO 0.05 45046
+ASVD 5.5 17632
+SliceGPT 0.6 18195
+GRASP 0.16 29050
+Table 11: Compression time and peak compression memory usage of GRASP and baselines for LLaMA2-7B under 25% compression ratio on a single A100 GPU.
+A.9 Robustness of GRASP towards different Representation Evaluation Dataset
+In this section, we provide details of the ablation studies conducted to investigate the impact of representation evaluation datasets and the amount of data used for singular value gradient attribution.
+Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of GRASP when compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
+Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of GRASP to the change in the number of calibration data.
 All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage.
-As shown in Figure 6, we can observe that Taco-SVD consistently achieves strong performance, indicating that our method is robust to variations in both the calibration dataset and the number of calibration data.
-Tables 8 and 9 summarize the results of Taco-SVD when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
-Figure 6: Performance of Taco-SVD on LLaMA3.1-8B-Instruct under 20% compression using (a) different calibration datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2.
-Taco-SVD demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
-Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
+As shown in Figure 5, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the representation evaluation dataset and the number of data.
+Tables 14 and 15 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different representation evaluation datasets (WikiText-2, C4) and varying numbers of calibration data.
+A.10 Robustness of GRASP towards Different Pruning Strategy
+In this section, we present the results of ablation experiments conducted with different pruning strategies in GRASP, with the results summarized in Table 12
+Pruning Strategy WikiText-2 PTB
+Average
+Accuracy
+Compression
+Time(h)
+One-shot Pruning 37.86 63.97 47.12 0.16
+Iterative Pruning 38.39 72.18 47.13 0.22
+Table 12: Comparison of one-shot pruning and iterative pruning for LLaMA3.1-8B-Instruct under 20% compression ratio.
+Model Summarization Few-shot Learning Synthetic Task
+1-1 1-2 1-3 1-4 1-1 2-1 2-2 2-3 2-4 3-1 3-2
+LLaMA3.1-8B-Instruct 28.35 20.04 25.85 14.51 63 51.3 39.29 16.5 1.64 9.67 17.17
+GRASP 25.76 19.41 25.97 8.99 59.5 67.44 38.41 18 1 10 18
+Model One-Doc QA Multi-Doc QA Code Completion Average
+4-1 4-2 4-3 4-4 5-1 5-2 5-3 5-4 6-1 6-2 ALL
+LLaMA3.1-8B-Instruct 9.96 6.04 18.75 14.27 11.27 11.57 7.65 34.16 56.47 51.34 26.09
+GRASP 15.32 20.47 34.52 16.46 31.33 26.18 10.28 29.86 33.74 38.09 26.26
+Table 13: Performance comparison of LLaMA3.1-8B-Instruct and its compressed version by GRASP under 20% compression ratio on LongBench.
+The datasets are grouped as follows: (1-1 to 1-4) denote GovReport, QMSum, MultiNews, and VCSUM; (2-1 to 2-4) denote TREC, TriviaQA, SAMSum, and LSHT; (3-1 to 3-3) denote PassageCount, PassageRetrieval-en, and PassageRetrieval-zh; (4-1 to 4-4) denote NarrativeQA, Qasper, MultiFieldQA-en, and MultiFieldQA-zh; (5-1 to 5-4) denote HotpotQA, 2WikiMultihopQA, MuSiQue, and DuReader; (6-1 to 6-2) denote LCC and RepoBench-P.
+Figure 5: Performance of GRASP on LLaMA3.1-8B-Instruct under 20% compression using (a) different representation evaluation datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2.
+GRASP demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
+Representation Evaluation Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
 WikiText-2 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
 C4 40.54 71.42 24 57.91 67.72 42.11 36.69 67.25 27.5 46.17
-Table 8: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by Taco-SVD under 20% compression using 512 samples from WikiText-2 and C4 as calibration datasets.
-Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
+Table 14: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression using 512 samples from WikiText-2 and C4 as representation evaluation datasets.
+Representation Evaluation Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
 WikiText-2-64 46.5 86.51 22.6 59.97 69.3 44.36 37.29 68.5 27.37 47.06
@@
 WikiText-2-512 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
-Table 9: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by Taco-SVD under 20% compression with varying calibration data sizes (64, 128, 256, 512) from WikiText-2.
+Table 15: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression with varying calibration data sizes (64, 128, 256, 512) from WikiText-2.

### v3

@@
-Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression
+GRASP: Replace Redundant Layers with Adaptive Singular Parameters for Efficient Model Compression
 Abstract
-Layer removal is an effective technique for compressing large language models (LLMs) by reducing redundancy and improving inference efficiency.
-However, indiscriminate pruning disrupts representation stability, leading to performance degradation.
-We propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), which preserves representation-critical singular values to mitigate these effects.
-Unlike direct layer removal, GRASP leverages gradient-based attribution on a syntax- and semantics-rich dataset to guide the selection of representation-critical singular values.
-By selectively applying singular value decomposition (SVD) to affected layers, GRASP achieves efficient compression while maintaining representation stability with minimal overhead.
-Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods in perplexity and downstream task performance.
+Recent studies have demonstrated that many layers are functionally redundant in large language models (LLMs), enabling model compression by removing these layers to reduce inference cost.
+While such approaches can improve efficiency, indiscriminate layer pruning often results in significant performance degradation.
+In this paper, we propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel compression framework that mitigates this issue by preserving sensitivity-aware singular values.
+Unlike direct layer pruning, GRASP leverages gradient-based attribution on a small calibration dataset to adaptively identify and retain critical singular components.
+By replacing redundant layers with only a minimal set of parameters, GRASP achieves efficient compression while maintaining strong performance with minimal overhead.
+Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods, achieving 90% of the original model’s performance under 20% compression ratio.
 The source code is available at GRASP.
@@
 […]
-Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression
+GRASP: Replace Redundant Layers with Adaptive Singular Parameters for Efficient Model Compression
 Kainan Liu1,2,†, Yong Zhang1,†, Ning Cheng1,*, Zhitao Li1, Shaojun Wang1, Jing Xiao1, 1Ping An Technology (Shenzhen) Co., Ltd., China 2The Hong Kong University of Science and Technology (Guangzhou) {zhangyong203, chengning211}@pingan.com.cn
@@
 To address this, model compression techniques such as quantization (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023), knowledge distillation (Gu et al., 2023; Xu et al., 2024), and pruning(Sun et al., 2024; Ashkboos et al., 2024) have been widely explored.
-Among these, structured pruning methods, especially layer removal (Men et al., 2024; Yang et al., 2024; Song et al., 2024), have shown promise by leveraging redundancy within layers of LLMs to achieve significant parameter reductions and inference speedup.
-Figure 1: Zero-shot performance of GRASP and pure layer removal under 20% compression ratio.
-Dashed lines represent perplexity on WikiText-2, while bar charts show the percentage of the original model’s average accuracy on seven reasoning datasets.
-Despite its simplicity and efficiency, layer removal introduces two main challenges, as shown in Figure 1.
-First, while task accuracy may remain stable in some cases, the perplexity (PPL) often degrades significantly, reflecting disruptions in the model’s internal representation consistency.
-This degradation arises from impaired information flow and misaligned intermediate representations (Liang et al., 2024).
-The sharp increase in PPL suggests that layer removal eliminates critical transformations necessary for maintaining coherent representation structures, potentially limiting the model’s generalization on complex or unseen tasks.
-Second, the impact of layer removal varies significantly across models.
-For example, under the same pruning ratio, LLaMA2 (Touvron et al., 2023b) maintains reasonable performance, while LLaMA3 (Dubey et al., 2024) experiences a steep performance drop.
-This discrepancy suggests that layer removal affects models unevenly, as certain structural transformations are more essential for maintaining stable representations, leading to performance inconsistencies across models.
-Our analysis suggests that these issues arise from the removal of transformations that play an essential role in preserving internal consistency and maintaining structured representations.
-In this paper, we introduce GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel hybrid pruning framework that integrates layer removal with gradient-based singular value selection for efficient LLM compression.
-Unlike conventional pruning techniques that either remove full layers or apply global SVD compression, GRASP selectively applies singular value decomposition (SVD) only to layers affected by pruning.
-A key component of GRASP is its use of a small, syntax- and semantics-rich dataset (e.g., WikiText-2), which serves as a representation evaluation dataset to guide the selection of singular values.
-Instead of relying solely on magnitude-based ranking, GRASP leverages gradient attribution computed from this dataset to dynamically identify and retain singular values critical for preserving language representations.
-This ensures that despite layer removal, the model retains essential syntactic and semantic structures, minimizing degradation in perplexity and downstream task performance.
-We evaluate GRASP on 10 datasets and five models from two distinct LLM families (LLaMA, Mistral).
-Extensive experiments validate the effectiveness of GRASP.
-It achieves compression rates comparable to pure layer removal while significantly improving perplexity and task performance across a wide rage of downstream tasks.
-Notably, GRASP maintains strong robustness across different models without requiring additional training.
-Furthermore, it achieves strong performance with minimal calibration data—using only 64 samples from a general-purpose, unlabeled dataset, demonstrating the efficiency of our approach.
-Additionally, post-training compensation further enhances GRASP’s performance, highlighting its adaptability in compressing large language models while preserving their linguistic capabilities.
-Figure 2: (A): Overview of GRASP.
-The figure illustrates how GRASP uses gradient attribution to identify representation-critical singular values and applies SVD to compress redundant layers while preserving representation-relevant transformations.
-(B): SVD-based LLM compression and (C): Layer removal compression methods
-2 Related Work
-2.1 Layer Removal-based LLM Compression
-Layer removal is a structured pruning technique that eliminates redundant layers within Large Language Models (LLMs), aiming to reduce the model’s computational and memory requirements.
-Methods such as ShortGPT (Men et al., 2024) introduce a metric called Block Influence to assess the significance of individual layers, enabling efficient one-shot removal of less important layers.
-SLEB (Song et al., 2024) improves this by employing an iterative pruning strategy, evaluating the importance of each layer based on the current state of the layer-removed LLMs.
-LaCo (Yang et al., 2024), on the other hand, proposes a gradual compression approach, progressively merging redundant layers from deeper to shallower parts of the network.
-Although effective in reducing model size, layer removal disrupts representation coherence, leading to performance degradation and increased perplexity, as analyzed by Liang et al.
-(2024).
-To mitigate this, post-training compensation methods have been proposed.
-Kim et al., 2024 introduced Shortened LLaMA, which employs LoRA (Hu et al., 2022) to restore pruned models’ capabilities.
-However, LoRA modifies the singular value spectrum, potentially weakening pre-trained features (Shuttleworth et al., 2024).
-LLM-Streamline (Chen et al., 2024) addresses this by training a lightweight module, such as an FFN or transformer layer, to approximate the removed layers.
-While effective, these methods impose high computational and data costs, limiting feasibility in resource-constrained settings.
-2.2 SVD-based LLM Compression
-Singular Value Decomposition (SVD) is widely used in model compression to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss.
-Recent methods have enhanced SVD to reduce LLM compression error.
-FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD.
-ASVD (Yuan et al., 2023) uses activation patterns from a representation evaluation dataset to scale weight matrices, reducing compression-induced activation errors.
-SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss.
-Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices.
-GRASP differs from these methods in three key aspects.
-First, rather than minimizing F2 norm reconstruction loss, GRASP optimizes a representation learning attribution loss, computed on a small representation evaluation dataset to ensure selection aligns with model representation rather than weight approximation.
-Second, instead of using a top-k singular value selection strategy, GRASP leverages gradient attribution to dynamically retain representation-critical singular values, which are not necessarily the largest in magnitude.
-Third, while SVD-based pruning is typically applied to all layers, GRASP is selectively used only in pruned layers, ensuring that layer removal does not disrupt critical transformations.
+Among these, structured pruning methods remove entire components such as neurons or layers to streamline the model, thereby achieving hardware efficiency and inference speedup.
+Figure 1: Unlike conventional layer pruning, which either skips redundant layers—often causing moderate performance drops—or replaces them with lightweight modules that require additional training, GRASP (right) retains only the most critical 10% of parameters within the redundant layers, effectively preserving accuracy with minimal overhead.
+In this work, we focus on structured layer pruning, which builds on prior findings that certain consecutive layers in large language models (LLMs) are functionally redundant.
+These findings have inspired approaches that either remove such layers entirely (Men et al., 2024, Yang et al., 2024, Kim et al., 2024) or replace them with lightweight modules (Chen et al., 2024) to reduce inference cost.
+Although layer removal is simple and computationally efficient, it often results in significant performance degradation.
+This degradation arises from disrupted information flow and misaligned intermediate representations (Liang et al., 2024), suggesting that layer removal eliminates certain important components that contribute meaningfully to maintaining model performance.
+Replacing redundant layers with lightweight modules can mitigate this issue to some extent, but such modules are typically randomly initialized, requiring substantial computational resources for training.
+In this paper, we propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel compression framework that replaces redundant layers with adaptive singular parameters for efficient LLM compression.
+Unlike direct layer removal, GRASP exploits the low-rank structure inherent in redundant layers, replacing the redundant layers with only a small subset of parameters while maintaining strong model performance.
+Specifically, GRASP operates in two key steps: First, it identifies layers suitable for pruning based on the cosine similarity of output hidden states between adjacent layers.
+Then, instead of relying on magnitude-based heuristics, GRASP leverages gradient attribution derived from a small calibration dataset to adaptively identify and retain the singular values most critical for downstream task performance.
+To evaluate the effectiveness of GRASP, We conduct extensive experiments on 19 datasets and 5 models from two distinct LLM families (LLaMA and Mistral).
+Notably, GRASP achieves strong performance in a training-free setting, requiring no additional optimization.
+Furthermore, when post-training compensation is applied, only a small number of samples are needed to rapidly restore model performance.
+This efficiency arises from retaining critical components within redundant layers, rather than relying on randomly initialized replacements.
+Overall, this paper makes the following contributions:
+  • We propose GRASP, a novel training-free compression framework that replaces redundant layers with adaptive singular parameters, leveraging the low-rank structure within LLMs to preserve performance with minimal overhead.
+  • We introduce a gradient-based singular value selection mechanism, enabling efficient identification of critical components without relying on magnitude-based heuristics.
+  • We conduct extensive experiments across ten datasets and five models from two major LLM families (LLaMA and Mistral), demonstrating that GRASP consistently achieves strong performance under both training-free and low-resource fine-tuning settings.
 2 Method
-3.1 Overview and Motivation
-In this paper, we propose GRASP, a novel hybrid gradient-guided compression framework designed to compress redundant layers in LLMs while preserving representation-critical transformations.
-Figure 2 provides an overview of GRASP.
-The method consists of two main steps: (1) identifying redundant layers based on cosine similarity, and (2) selecting and retaining representation-critical singular values and their corresponding singular vectors using gradient-based attribution.
+Figure 1 illustrates the workflow of GRASP.
+The method consists of two main steps: Identifying redundant layers to be pruned (Sec 2.1) and replacing redundant layers with critical singular components guided by gradient-based attribution (Sec 2.2).
 Below, we describe each step in detail.
@@
 Instead of directly removing these layers, GRASP compresses weight matrices with a gradient-based approach to retain critical internal transformations.
-3.3 Gradient-Guided Singular Value Selection
-3.3.1 Representation Learning Attribution Loss (RLAL)
+2.2 Layer Replacement with Adaptive Singular Parameters
+Motivation.
+GRASP operates under the hypothesis that redundant layers exhibit an inherent low-rank structure, allowing their functionality to be effectively approximated using low-rank matrices.
+Based on this insight, a straightforward approach is to apply singular value decomposition (SVD) to these layers.
+However, prioritizing components solely based on singular value magnitude does not necessarily correlate with downstream task performance (Hsu et al., 2022; Hua et al., 2025).
+To validate this point, we selectively zero out groups of singular values in the weight matrices of large language models and measure their impact on downstream tasks.
+As shown in Figure 2, we make two key observations: (1) The contribution of a singular value to downstream task performance is not solely determined by its magnitude; smaller values can be crucial for task performance in some cases.
+(2) Redundant layers exhibit a highly low-rank structure, where only a few singular directions dominate the model performance.
+Figure 2: Sensitivity analysis of grouped singular value truncation. While singular values are typically ordered by magnitude, their impact on downstream performance does not follow the same order.
+Key Design.
 To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value based on its contribution to model performance, rather than relying on magnitude alone.
-Unlike traditional SVD-based methods that select top-$k$ singular values using a reconstruction loss measured by the Frobenius norm ($F_{2}$ norm), GRASP utilizes a Representation Learning Attribution Loss (RLAL), designed specifically for general representation capability modeling.
-RLAL is computed on a small, unlabeled representation evaluation dataset derived from corpora rich in syntactic and semantic structures, such as WikiText-2 and similar datasets.
-These datasets ensure that the retained singular values are those most crucial for maintaining the model’s ability to capture fundamental language structures, including syntax and semantics, even after compression.
-Unlike task-specific objectives, RLAL focuses on preserving linguistic consistency and general-purpose representation learning rather than optimizing for any particular downstream application.
-Mathematically, RLAL is formulated as:
-$\mathcal{L}_{\text{RLAL}}=-\sum_{t}\log P(y_{t}\mid x_{\leq t})$ (2)
-where $y_{t}$ is the target token at time step $t$, $x_{\leq t}$ denotes the sequence of preceding tokens, and $P(y_{t}\mid x_{\leq t})$ represents the predicted probability of the next token given the context.
-This loss function aligns with the standard language modeling loss, ensuring that the model learns meaningful representations.
-3.3.2 Singular Value Decomposition (SVD)
-For a redundant layer’s weight matrix $W_{\text{redundant}}\in\mathbb{R}^{m\times n}$, SVD decomposes $W_{\text{redundant}}$ as:
-$W_{\text{redundant}}=U\Sigma V^{T}$ (3)
-where $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{k})$ contains the singular values ($\sigma_{i}\geq 0$), and $U\in\mathbb{R}^{m\times m}$, $V\in\mathbb{R}^{n\times n}$ are orthogonal matrices containing the left and right singular vectors, respectively.
-3.3.3 Representation-Critical Singular Value Importance
-To assess the representation-critical importance of each singular value, we introduce $\Phi(\sigma_{i})$, the Representation-Critical Singular Value Importance, which quantifies the contribution of $\sigma_{i}$ to the $\mathcal{L}_{\text{RLAL}}$.
-Using its first-order Taylor expansion, the alteration in the $\mathcal{L}_{\text{RLAL}}$ caused by $\sigma_{i}$ is approximated as:
-$\Delta\mathcal{L}_{\text{RLAL}}(\sigma_{i})\approx\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}\Delta\sigma_{i}$ (4)
-where $\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}$ represents the gradient sensitivity of $\sigma_{i}$ to the representation learning attribution loss , and $\Delta\sigma_{i}$ is the perturbation applied to $\sigma_{i}$.
-The gradient sensitivity is calculated by projecting the gradient of $\mathcal{L}_{\text{RLAL}}$ with respect to $W_{\text{redundant}}$ onto the singular directions:
-$\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}=\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$ (5)
-where $G=\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial W}$ is the gradient of the $\mathcal{L}_{\text{RLAL}}$ with respect to $W_{\text{redundant}}$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the left and right singular vectors corresponding to $\sigma_{i}$.
-The detailed proof can be found in the Appendix A.1.
-Integrating the gradient sensitivity with the magnitude of $\sigma_{i}$, we define $\Phi(\sigma_{i})$ as:
-$\Phi(\sigma_{i})=\sigma_{i}\cdot\left(\mathbf{u}_{i}^{T}G\mathbf{v}_{i}\right)$ (6)
-where $\Phi(\sigma_{i})$ measures the overall contribution of $\sigma_{i}$ to representation performance, considering both its magnitude and alignment with the gradient of the general-purpose model loss.
-A larger value of $\Phi(\sigma_{i})$ indicates greater importance.
-We rank singular values in descending order of representation relevance based on $\Phi(\sigma_{i})$.
-Only the top-$k$ singular values with the highest $\Phi(\sigma_{i})$ are retained, and the compressed weight matrix is reconstructed as:
-$W_{\text{RLAL}}=\sum_{i=1}^{k}\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{T}$ (7)
-This ensures that $W_{\text{RLAL}}$ preserves the transformations most critical to representation stability, achieving efficient compression while maintaining representation consistency.
-3.4 Layer-wise Low-Rank Decomposition
-GRASP applies a layer-wise backward compression strategy, starting from the last redundant layer and moving sequentially backward.
-For each redundant layer, weight matrices are compressed using Gradient-Guided Singular Value Decomposition, retaining only the top-$k$ singular values ranked by $\Phi(\sigma_{i})$.
-This ensures a low-rank approximation that preserves representation-critical transformations.
-By processing one layer at a time, this approach significantly reduces memory overhead, avoiding the need to load the entire model into memory.
-Additionally, the backward order ensures that compressing deeper layers does not disrupt the consistency of earlier layers, preserving the overall stability of the model.
-The detailed implementation for Layer-wise Low-Rank Decomposition is provided in Algorithm 1 in the Appendix A.2.
+Formally, the singular value decomposition of a weight matrix $W\in\mathbb{R}^{m\times n}$ is given by $W=U\Sigma V^{\top}$.
+Equivalently, $W$ can be represented as the sum of its rank-one components:
+$W=\sum_{k=1}^{l}u_{k}\sigma_{k}v_{k}^{\top},$ (2)
+where $l=\min(m,n)$, $\sigma_{k}$ denotes the $k$-th singular value, and $u_{k}$, $v_{k}$ are the $k$-th column vectors of the orthogonal matrices $U$ and $V$, respectively.
+For brevity, we use $\Phi_{k}$ to represent the singular group $\{u_{k},\sigma_{k},v_{k}^{\top}\}$.
+GRASP estimates the importance of each $\Phi_{k}$ using a small, general-purpose calibration dataset (e.g., WikiText-2), computing a sensitivity-based score that reflects its effect on model performance, given by:
+$I(\Phi_{k})=T(\sigma_{k})+\sum_{i=1}^{m}T(u_{k,i})+\sum_{i=1}^{n}T(v_{k,i})$ (3)
+where $T(\cdot)$ denotes the estimated change in loss when a certain parameter $\theta$ is zeroed out.
+This can be approximated by the second-order Taylor series expansion (LeCun et al., 1989), which is defined as:
+$T(\theta)=\left|\theta^{\top}\nabla_{\theta}\mathcal{L}+\frac{1}{2}\theta^{\top}H\theta+\mathcal{O}(\|\theta\|^{3})\right|$ (4)
+$\mathcal{L}=-\sum_{t}\log P(y_{t}\mid x_{\leq t})$ (5)
+where $\nabla_{\theta}\mathcal{L}$ denotes the gradient of the standard language modeling objective function $\mathcal{L}$ with respect to the parameter $\theta$, and $H$ represents the corresponding Hessian matrix.
+To reduce computational overhead, we omit the second-order term and approximate the importance using only the first-order derivative (Hua et al., 2025, Kim et al., 2024):
+$I(\Phi_{k})=\left|\sigma_{k}\frac{\partial\mathcal{L}}{\partial\sigma_{k}}\right|+\sum_{i=1}^{m}\left|u_{k,i}\frac{\partial\mathcal{L}}{\partial u_{k,i}}\right|+\sum_{i=1}^{n}\left|v_{k,i}\frac{\partial\mathcal{L}}{\partial v_{k,i}}\right|$ (6)
+The first term in Eq.
+6 captures the gradient projection of the loss $\mathcal{L}$ with respect to the singular value $\sigma_{i}$.
+The detailed proof is provided in Appendix A.1.
+By aggregating the first-order expansions across all components within each singular group, our method effectively captures the contribution of singular values.
+A larger value indicates a greater importance of the singular group.
+This gradient-oriented attribution moves beyond heuristic magnitude-based criteria, enabling performance-aligned importance evaluation.
+Under the hypothesis that redundant layers exhibit an inherent low-rank structure, we retain only the top-$k$ singular values most critical to model performance based on Eq.
+6, and use them to replace the corresponding redundant layer.
+2.3 Detailed Implementation of GRASP
+GRASP processes the redundant layers in a sequential manner, starting from the final redundant layer and proceeding backward through the network. We summarize the detailed algorithm in Algorithm 1.
+Algorithm 1 GRASP: Gradient-based Retention of Adaptive Singular Parameters
+1:Model $M$, Calibration set $D$, Retain ratio $r$
+2:Compressed model $\tilde{M}$
+3:Step1: Redundant Layer Selection:
+4:Compute cosine similarity $\text{cos}(H_{i},H_{i+1})$ for all layers
+5:Select top-$L$ layers with highest similarity as redundant
+6:Step 2: Gradient-Guided Compression:
+7:for each redundant layer $l$ (in reverse order) do
+[…]
+9:         SVD: $W=U\Sigma V^{\top}$
+10:         Compute importance $I(\Phi_{k})$ for each $\Phi_{k}=\{u_{k},\sigma_{k},v_{k}^{\top}\}$ via Eq. 6
+11:         Keep top-$r\%$ singular groups, reconstruct $\tilde{W}$
+12:     end for
+13:end for
+14:return $\tilde{M}$
 3 Experiments
-In this section, we conduct comprehensive experiments to evaluate GRASP from four key perspectives.
-(1) We first compare our method with several popular pruning-based LLM compression methods to demonstrate its effectiveness (Section 4.2).
-(2) Next, we compare GRASP with SVD-based LLM compression methods under different compression ratios with different LLMs (Section 4.3).
-(3) Then, we further evaluate GRASP on the LongBench (Bai et al., 2023) datasets to assess its performance on long-form text processing tasks (Section 4.4).
-(4) Finally, we analyze the inference speed-up achieved by GRASP (Section 4.5).
+In this section, we conduct comprehensive experiments to evaluate GRASP from three key perspectives.
+(1) We first compare our method with existing pruning-based LLM compression approaches to demonstrate its effectiveness (Section 3.2).
+(2) Next, we analyze the inference speed-up achieved by GRASP (Section 3.3).
+(3) Finally, we investigate the factors influencing our approach by performing ablation studies on the choice of calibration datasets and pruning strategies.
+(Section 3.4)
 3.1 Experimental Setup
-Below we detail the models, benchmarks and implementation details, with more experimental setups provided in Appendix A.3 due to page limit.
-Models
-We evaluated our method on a diverse set of LLMs including LLaMA family models LLaMA-7B (Touvron et al., 2023a), LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023b), LLaMA3.1-8B-Instruct (Dubey et al., 2024) and Mistral family models Mistral-7B (Jiang et al., 2023).
-Benchmark
-10 datasets are used as evaluation benchmarks including two pre-training datasets: WikiText-2 (Merity et al., 2017), PTB (Marcus et al., 1993), seven common sense reasoning datasets include: WinoGrande (Sakaguchi et al., 2020), Hellaswag (Zellers et al., 2019), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-c, ARC-e (Clark et al., 2018), MathQA (Amini et al., 2019) and one generation benchmark GSM8K (Cobbe et al., 2021).
-Additionally, we include LongBench (Bai et al., 2023) to assess LLM’s performance on long-form text processing after pruning.
-Implementation Details
-For all baselines, we randomly sample 512 instances from the WikiText-2 dataset as the representation evaluation dataset.
-For pruning-based with post-training compensation baselines, all models are trained on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32.
+Below we detail the models, benchmarks, baselines and implementation details used in our experiments, with more experimental setups provided in Appendix A.2.
+Models.
+We evaluate GRASP on a range of large language models (LLMs) from two model families: the LLaMA family, including LLaMA-7B (Touvron et al., 2023a), LLaMA 2-7B, LLaMA 2-13B (Touvron et al., 2023b), and LLaMA 3.1-8B-Instruct (Dubey et al., 2024), as well as Mistral-7B (Jiang et al., 2023) from the Mistral family.
+Baselines.
+We compare GRASP against 8 structured pruning methods to substantiate its efficacy:
+  • Layer-pruning methods We consider three representative layer-pruning methods as baselines: ShortGPT (Men et al., 2024), LaCo (Yang et al., 2024) and LLM-Streamline (Chen et al., 2024).
+  • Module-pruning methods We also select LLM-Pruner (Ma et al., 2023) and SliceGPT (Ashkboos et al., 2024) which prune the redundant modules in LLMs.
+  • Low-rank Pruning methods Considering our method involves Gradient-based SVD, we also compare with other low-rank pruning methods: FWSVD (Hsu et al., 2022), ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024).
+We provide a detailed comparison of these pruning-based LLM compression methods in Appendix A.3.
+Implementation Details.
+To ensure a fair comparison, we randomly sample 512 data points from the WikiText-2 dataset as the calibration dataset.
+All experiments are conducted on NVIDIA A100-SXM4 (80GB) GPUs.
 Further experimental details can be found in Appendix A.2.
-Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
-Method
-PPL $\downarrow$
-(WikiText-2)
-PPL $\downarrow$
-Acc $\uparrow$
-Average
-PPL $\downarrow$
-(WikiText-2)
-PPL $\downarrow$
-Acc $\uparrow$
-Average
-PPL $\downarrow$
-(WikiText-2)
-PPL $\downarrow$
-Acc $\uparrow$
-Average
-PPL $\downarrow$
-(WikiText-2)
-PPL $\downarrow$
-Acc $\uparrow$
-Average
-Original 5.25 23.73 0.59 5.47 20.83 0.53 4.88 28.92 0.56 7.21 12.33 0.60
-ShortGPT 23.95 80.70 0.46 18.45 57.58 0.43 27.26 59.14 0.44 140.47 262.47 0.44
-LaCo 244.53 524.26 0.34 39.43 93.98 0.43 38.27 125.98 0.42 753.98 1577.44 0.43
-SliceGPT 9.04 107.65 0.39 8.63 99.25 0.41 7.44 99.77 0.45 52.78 460.73 0.35
-Ours 18.42 77.98 0.46 16.12 44.09 0.44 11.1 56.57 0.47 37.86 63.97 0.47
-Table 1: Perplexity($\downarrow$) of GRASP and pruning-based without post-training compensation baselines on the WikiText-2 and PTB datasets, as well as the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 20% compression ratio.
-The best performance is marked in bold.
-Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
-Method
-PPL $\downarrow$
-(WikiText-2)
-PPL $\downarrow$
-Acc $\uparrow$
-Average
-PPL $\downarrow$
-(WikiText-2)
-PPL $\downarrow$
-Acc $\uparrow$
-Average
-PPL $\downarrow$
-(WikiText-2)
-PPL $\downarrow$
-Acc $\uparrow$
-Average
-PPL $\downarrow$
-(WikiText-2)
-PPL $\downarrow$
-Acc $\uparrow$
-Average
-Original 5.25 23.73 0.59 5.47 20.83 0.53 4.88 28.92 0.56 7.21 12.33 0.60
-LoRA 11.56 34.05 0.51 11.7 35.39 0.46 8.46 58.15 0.50 14.56 21.24 0.52
-LS-FFN 16.12 59.98 0.49 13.94 37.43 0.46 8.11 76.3 0.50 16.79 24.76 0.51
-LS-Layer 14.67 54.31 0.49 11.57 32.62 0.47 8.23 77.04 0.50 14.27 20.85 0.53
-Ours* 11.62 44.14 0.51 9.59 30.38 0.48 7.52 38.31 0.51 14.13 20.99 0.53
-Table 2: Perplexity($\downarrow$) of GRASP (Ours* denotes the version with post-training compensation) and pruning-based with post-training compensation baselines on the WikiText-2 and PTB datasets, as well as the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 25% compression ratio.
-LS-FFN and LS-Layer denote two variants of LLM-Streamline.
-The best performance is marked in bold.
-Ratio Method WikiText-2 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA GSM8K Average Percentage
-0% Original 5.68 0.28 0.67 0.67 0.56 0.38 0.78 0.27 0.09 0.46 100.0
-20% FWSVD 1727 0.15 0.31 0.50 0.26 0.23 0.56 0.21 0.00 0.28 60.8
-ASVD 11.14 0.25 0.53 0.64 0.41 0.27 0.68 0.24 0.04 0.38 82.6
-SVD-LLM 7.94 0.22 0.58 0.63 0.43 0.29 0.69 0.24 0.05 0.39 84.7
-Ours 14.79 0.22 0.52 0.64 0.43 0.32 0.70 0.24 0.03 0.39 84.7
-Ours* 10.19 0.24 0.59 0.63 0.5 0.35 0.73 0.25 0.04 0.42 91.3
-30% FWSVD 20127 0.17 0.26 0.49 0.26 0.22 0.51 0.19 0.00 0.26 56.5
-ASVD 51 0.18 0.43 0.53 0.37 0.25 0.65 0.21 0.00 0.33 71.7
-SVD-LLM 9.56 0.20 0.48 0.59 0.37 0.26 0.65 0.22 0.03 0.35 76.1
-Ours 30.48 0.19 0.42 0.62 0.39 0.28 0.64 0.23 0.02 0.35 76.1
-Ours* 12.63 0.24 0.54 0.64 0.46 0.32 0.69 0.24 0.04 0.40 87.0
-40% FWSVD 18156 0.16 0.26 0.51 0.26 0.22 0.53 0.21 0.00 0.27 58.7
-ASVD 15358 0.13 0.28 0.48 0.26 0.22 0.55 0.19 0.00 0.26 56.5
-SVD-LLM 13.11 0.19 0.42 0.58 0.33 0.25 0.60 0.21 0.02 0.33 71.7
-Ours 62.34 0.18 0.37 0.57 0.35 0.27 0.61 0.21 0.01 0.32 69.6
-Ours* 16.25 0.22 0.49 0.63 0.43 0.3 0.68 0.23 0.02 0.38 82.6
-50% FWSVD 24391 0.12 0.26 0.50 0.26 0.23 0.53 0.20 0.00 0.26 56.5
-ASVD 15358 0.12 0.26 0.51 0.26 0.22 0.52 0.19 0.00 0.26 56.5
-SVD-LLM 23.97 0.16 0.33 0.54 0.29 0.23 0.56 0.21 0.00 0.29 63.0
-Ours 102.58 0.13 0.29 0.53 0.28 0.23 0.53 0.20 0.01 0.28 60.9
-Ours* 23.69 0.18 0.4 0.56 0.35 0.26 0.61 0.21 0.02 0.32 69.6
+Methods OpenbookQA ARC_e WinoGrande HellaSwag ARC_c PIQA MathQA Average Percentage
+Dense 0.34 0.82 0.74 0.59 0.52 0.80 0.39 0.60 100.0%
+LaCo 0.26 0.49 0.65 0.33 0.30 0.65 0.30 0.42 70.9%
+ShortGPT 0.21 0.57 0.66 0.42 0.32 0.67 0.26 0.44 74.1%
+SliceGPT 0.15 0.43 0.51 0.30 0.23 0.58 0.22 0.35 57.7%
+GRASP 0.22 0.60 0.70 0.44 0.37 0.69 0.28 0.47 78.6%
+Table 1: Zero-shot performance of GRASP and structured pruning baselines without post-training under a 20% compression ratio.
+Results are reported on seven reasoning datasets (individual and average accuracy).
+Bold values indicate the best performance.
+Method C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU Avg. Per.
+Dense 43.8 33.0 41.6 70.8 37.5 66.7 71.3 78.1 33.1 35.5 46.8 31.8 49.2 100.0%
+LLMPruner* 29.7 33.4 28.4 58.7 40.4 48.5 54.6 72.0 22.9 22 25.3 25.0 38.4 78.1%
+SliceGPT* 31.5 31.6 18.5 59.9 43.3 49.6 47.5 68.3 27.0 29.4 28.8 24.8 38.4 78.0%
+LaCo* 39.7 34.4 36.1 64.1 40.4 45.7 55.7 69.8 23.6 22.6 26.5 25.2 40.3 82.0%
+ShortGPT* 40.2 34.4 21.5 67.3 40.4 51.7 59.7 69.0 35.2 34.7 44.6 28.9 44.0 89.4%
+LLM-Streamline-FFN* 40.7 33.0 22.8 65.9 38.5 60.6 61.2 71.2 38.0 38.7 47.0 31.7 45.8 93.1%
+LLM-Streamline-Layer* 43.3 33.0 24.1 67.5 36.5 59.2 61.1 71.5 34.8 37.0 45.5 29.4 45.2 92.0%
+GRASP 44.6 35.1 26.2 68.4 41.4 63.2 62.7 73.3 35.1 36.1 43.1 30.7 46.7 94.9%
+Table 2: Comparison between GRASP and structured pruning baselines with post-training compensation under a 20% compression ratio.
+Results marked with * are reported from (Chen et al., 2024).
+Bold values indicate the best performance.
+Ratio Method Openb. ARC_e WinoG. HeSW ARC_c PIQA MathQA GSM8K Average Percentage
+0% Original 0.28 0.67 0.67 0.56 0.38 0.78 0.27 0.09 0.46 100.0
+20% FWSVD 0.15 0.31 0.50 0.26 0.23 0.56 0.21 0.00 0.28 60.8
+ASVD 0.25 0.53 0.64 0.41 0.27 0.68 0.24 0.04 0.38 82.6
+SVD-LLM 0.22 0.58 0.63 0.43 0.29 0.69 0.24 0.05 0.39 84.7
+Ours 0.22 0.52 0.64 0.43 0.32 0.70 0.24 0.03 0.39 84.7
+Ours* 0.24 0.59 0.63 0.5 0.35 0.73 0.25 0.04 0.42 91.3
+30% FWSVD 0.17 0.26 0.49 0.26 0.22 0.51 0.19 0.00 0.26 56.5
+ASVD 0.18 0.43 0.53 0.37 0.25 0.65 0.21 0.00 0.33 71.7
+SVD-LLM 0.20 0.48 0.59 0.37 0.26 0.65 0.22 0.03 0.35 76.1
+Ours 0.19 0.42 0.62 0.39 0.28 0.64 0.23 0.02 0.35 76.1
+Ours* 0.24 0.54 0.64 0.46 0.32 0.69 0.24 0.04 0.40 87.0
+40% FWSVD 0.16 0.26 0.51 0.26 0.22 0.53 0.21 0.00 0.27 58.7
+ASVD 0.13 0.28 0.48 0.26 0.22 0.55 0.19 0.00 0.26 56.5
+SVD-LLM 0.19 0.42 0.58 0.33 0.25 0.60 0.21 0.02 0.33 71.7
+Ours 0.18 0.37 0.57 0.35 0.27 0.61 0.21 0.01 0.32 69.6
+Ours* 0.22 0.49 0.63 0.43 0.3 0.68 0.23 0.02 0.38 82.6
+50% FWSVD 0.12 0.26 0.50 0.26 0.23 0.53 0.20 0.00 0.26 56.5
+ASVD 0.12 0.26 0.51 0.26 0.22 0.52 0.19 0.00 0.26 56.5
+SVD-LLM 0.16 0.33 0.54 0.29 0.23 0.56 0.21 0.00 0.29 63.0
+Ours 0.13 0.29 0.53 0.28 0.23 0.53 0.20 0.01 0.28 60.9
+Ours* 0.18 0.4 0.56 0.35 0.26 0.61 0.21 0.02 0.32 69.6
 […]
-Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
-Method
-PPL $\downarrow$
-(WikiText-2)
-Acc $\uparrow$
-Average
-PPL $\downarrow$
-(WikiText-2)
-Acc $\uparrow$
-Average
-PPL $\downarrow$
-(WikiText-2)
-Acc $\uparrow$
-Average
-PPL $\downarrow$
-(WikiText-2)
-Acc $\uparrow$
-Average
-Original 5.25 0.59 5.68 0.52 5.47 0.53 7.21 0.60
-FWSVD 6357 0.32 1727 0.32 2360 0.31 3256.7 0.29
-ASVD 19.28 0.4 11.14 0.44 9.70 0.46 2443.99 0.30
-SVD-LLM 10.21 0.41 7.94 0.44 8.50 0.47 - -
-Ours 18.42 0.45 14.79 0.44 16.12 0.44 37.86 0.47
-Ours* 11.62 0.51 10.19 0.47 9.59 0.48 14.13 0.53
-Table 4: Perplexity($\downarrow$) of GRASP and SVD-based baselines on the WikiText-2 datasets and the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 20% compression ratio.
-"†" indicates that we refer to the results in the original paper.
-The best performance is marked in bold.
-Model
-Single-Doc
-QA
-Multi-Doc
-QA
-Summarization
-Task
-Few-shot
-Learning
-Synthetic
-Task
-Code
-Completion
-Average
-LLaMA3.1-8B-Instruct 12.26 16.16 22.19 42.52 9.49 53.91 26.09
-GRASP* 21.69 24.41 20.03 45.84 9.67 35.92 26.26
-Table 5: Evaluations of LLaMA3.1-8B-Instruct and its compressed version of GRASP* under 20% compression ratio on LongBench benchmarks.
-The average results across six task categories are shown in this table, with additional details provided in the Appendix A.7.
 3.2 Comparison with Pruning-based LLM Compression Methods
-Post-training compensation plays a crucial role in restoring the performance of compressed LLMs.
-To ensure a fair comparison, we compare GRASP with two categories of pruning-based compression methods.
-(1) Pruning without post-training compensation, including two layer removal methods, ShortGPT (Men et al., 2024), LaCo (Yang et al., 2024), and state-of-the-art structured pruning method: SliceGPT (Ashkboos et al., 2024).
-(2) Pruning with post-training compensation: ShortGPT with LoRA (Hu et al., 2022) and LLM-Streamline (Chen et al., 2024).
-We provide a detailed comparison of these pruning-based LLM compression methods in Appendix A.4.
-Pruning without post-training compensation
-We present the perplexity on WikiText-2 and PTB datasets, as well as the average accuracy on common sense reasoning benchmarks for different methods in Table 1.
-The results demonstrate that GRASP consistently achieves strong performance across all models and tasks.
-Specifically, in common sense reasoning tasks, GRASP surpasses SliceGPT by 15% in accuracy on Mistral-7B and by 34% on LLaMA3.1-8B-Instruct.
-For language generation tasks, while SliceGPT achieves lower perplexity on the WikiText-2 dataset, GRASP consistently outperforms the baselines on the PTB dataset.
-We hypothesize this discrepancy arises from the potential overfitting problem of baselines to the calibration dataset.
-Notably, GRASP exhibits significantly improved robustness compared to baselines, effectively addressing the variability in sensitivity to pruning across diverse models.
-Detailed results for each common sense reasoning benchmark are provided in Appendix A.5.
-Pruning with Post-Training Compensation
-We further demonstrate that GRASP provides a better initialization for post-training compensation, enabling more accurate recovery of the compressed model compared to other methods.
+3.2.1 Experiment 1
+Models and Benchmarks.
+In this experiment, we evaluate GRASP against representative structured pruning baselines using a more modern LLM architecture—LLaMA 3.1-8B-Instruct—without applying any post-training compensation.
+The model is compressed to 20% of its original size and evaluated on seven commonsense reasoning benchmarks, including WinoGrande (Sakaguchi et al., 2020), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-e, ARC-c (Clark et al., 2018), and MathQA (Amini et al., 2019).
+All tasks are tested in a zero-shot setting using the LM-Evaluation-Harness framework (Gao et al., 2024).
+Main Results.
+As shown in Table 1, GRASP achieves the highest average accuracy across seven commonsense reasoning benchmarks, consistently outperforming all baseline methods.
+In particular, GRASP improves over SliceGPT by 34% in average accuracy and outperforms LaCo by 12%.
+To further evaluate the generalizability of GRASP across different LLM architectures, we additionally conduct experiments on LLaMA 2-7B, LLaMA 2-13B, and Mistral-7B.
+Detailed results are presented in Appendix A.4.
+Notably, GRASP demonstrates significantly improved robustness across models, effectively mitigating the variability in pruning sensitivity across diverse architectures.
+3.2.2 Experiment 2
+Models and Benchmarks.
+In this section, following prior research, we compress the LLaMA 2-7B model under a 25% compression ratio and evaluate the compressed model on a broad set of natural language understanding (NLU) and question-answering (QA) benchmarks, including CMNLI (Xu et al., 2020), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), CHID (Zheng et al., 2019), WSC (Levesque et al., 2012), CommonsenseQA (Talmor et al., 2018), BoolQ (Clark et al., 2019), MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2023), Race (Lai et al., 2017) and C3 (Sun et al., 2020).
+We utilized the OpenCompass evaluation framework (Contributors, 2023) and report accuracy as the evaluation metric for all benchmarks under the PPL mode, following the same evaluation protocol as LaCo (Yang et al., 2024).
+Main Results.
 To ensure a fair comparison, we constrain the number of trainable parameters to remain approximately the same across all methods by retaining only 10% of the parameters in each redundant layer and allowing only these parameters to be trainable.
-As shown in Table 2, GRASP consistently outperforms LLM-Streamline across different LLMs.
-Furthermore, we present the validation loss curves for each method during post-training in Appendix A.6, which show that GRASP GRASP enables more stable and faster convergence.
-More Detailed results are provided in Appendix A.5.
-4.3 Comparison with SVD-based LLM Compression Methods
-In this section, we compare the performance of GRASP with state-of-the-art SVD-based LLM compression methods—FWSVD (Hsu et al., 2022), ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024)—from two distinct perspectives: (1) performance under different compression ratios and (2) performance across different LLMs.
-Performance under Different Compression Ratios.
-To evaluate the performance and stability of our proposed method under various compression ratios, we compare GRASP with the SVD-based baselines on LLaMA-7B under compression ratios ranging from 20% to 50%.
+Notably, for GRASP’s post-training compensation process, we fine-tune the compressed model on Alpaca (Taori et al., 2023) for only one epoch to guarantee efficiency.
+Additional implementation details are provided in Appendix A.2.
+As shown in Table 2, GRASP consistently outperforms the best-performing baselines LLM-Streamline on average and achieves a 94.9% of the original model performance at a 25% compression ratio.
+In addition to achieving superior accuracy, GRASP also demonstrates more stable and faster convergence during post-training, which we attribute to its preservation of critical singular components.
+3.2.3 Experiment 3
+Models and Benchmarks.
+In this section, we further compare our method against state-of-the-art structured low-rank pruning approaches—FWSVD (Hsu et al., 2022), ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024) on the LLaMA-7B model under various compression ratio.
+8 datasets are used as evaluation benchmarks including seven commonsense reasoning datasets as Experiment 1 and one natural language generation (NLG) benchmark GSM8K (Cobbe et al., 2021).
+For all benchmarks, we report the zero-shot accuracy as the evaluation metric.
+Main Results.
+To evaluate the performance and stability of our proposed method, we conduct experiments under various compression ratios ranging from 20% to 50%.
 Table 3 summarizes the results for different methods.
-The results demonstrate that our proposed GRASP consistently outperforms the baseline methods on most benchmarks, except for a slightly higher perplexity on WikiText-2 compared to SVD-LLM.
+The results demonstrate that our proposed GRASP consistently outperforms the baseline methods on most benchmarks.
 Specifically, GRASP retains more than 91% of the original performance at a 20% compression ratio and 87% under the compression ratio of 30%.
-More importantly, with fast and resource-efficient post-training, GRASP enables rapid accuracy recovery, achieving 70% of the original model performance even at a 50% compression ratio.
-Performance on different LLMs.
-To evaluate the generalizability of GRASP across different LLMs, we compare its performance with the SVD-based baselines under 20% compression ratio on four models from two distinct LLM families-LLaMA 2-7B, LLaMA 2-13B, LLaMA 3.1-8B-Instruct and Mistral-7B.
-The results, presented in Table 4, indicate that while existing methods achieve competitive performance on the LLaMA 2 series, they exhibit a sharp performance drop on other architectures, particularly on LLaMA 3.
-In contrast, GRASP consistently outperforms baseline methods across all architectures and demonstrates greater stability across different LLMs.
-4.4 Performance in Long-Form Text Processing
-In this section, we evaluate the performance of GRASP on LongBench (Bai et al., 2023) to assess its performance on long-form reasoning and complex generative tasks.
-The average performance across six task categories: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion are summarized in Table 5, with detailed results provided in the Appendix A.7.
-As shown, the compressed version by GRASP under 20% compression ratio still maintains strong performance and even surpasses the original LLaMA3.1-8B-Instruct model in single-document QA, multi-document QA, and few-shot learning tasks.
-This demonstrates the effectiveness of GRASP in rapidly recovering accuracy and generative capability through post-training compensation after pruning.
+More importantly, with fast and resource-efficient post-training compensation, GRASP enables rapid accuracy recovery, achieving 70% of the original model performance even at a 50% compression ratio.
+Furthermore, to assess the generalizability of GRASP across different LLM architectures, we compare its performance against structured low-rank pruning baselines under a 20% compression ratio on four models from two distinct LLM families: LLaMA 2-7B, LLaMA 2-13B, LLaMA 3.1-8B-Instruct, and Mistral-7B.
+As shown in Figure 3, GRASP consistently outperforms all baseline methods across architectures and exhibits greater robustness across different model families.
+The only exception is on LLaMA 2-13B, where GRASP slightly underperforms SVD-LLM.
+However, this performance gap can be quickly recovered through lightweight compensation.
+The detailed results are provided in Appendix A.5.
+Figure 3: Comparison between our method and low-rank pruning baselines on four different LLMs.
+Average accuracy is reported across seven commonsense reasoning benchmarks: OpenBookQA, WinoGrande, HellaSwag, ARC-easy, ARC-challenge, PIQA, and MathQA.
+Figure 4: Performance of GRASP on LLaMA3.1-8B-Instruct under 20% compression using (a) different calibration datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2.
+GRASP demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
 3.3 Compression Costs and Inference Efficiency of GRASP
-GRASP can compress LLMs at very low cost while improve the inference efficiency on real hardware.
-To evaluate its acceleration effect, we compared the throughput (tokens generated per second) of original LLaMA2-7B and its compressed version by GRASP with different sequence lengths and batch sizes.
-As shown in Figure 3, GRASP consistently consistently improves generation speed and achieves an acceleration effect comparable to pure layer removal.
-It is important to note that although GRASP retains a small portion of parameters in the redundant layers to mitigate the internal inconsistencies introduced by layer removal, this structure is extremely low-rank, resulting in negligible inference latency while maintaining task performance.
-Additionally, we also measure the compression time and peak memory consumption of GRASP when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio.
-The detailed results are provided in Appendix A.8.
+GRASP enables low-cost compression of LLMs while improving inference efficiency on real hardware.
+To evaluate its acceleration benefits, we measure the throughput (tokens per second) of the original LLaMA2-7B and its GRASP-compressed counterpart under varying sequence lengths and batch sizes.
+As shown in Figure 5, GRASP consistently improves generation speed and achieves acceleration comparable to direct layer removal.
+Notably, although GRASP retains a small subset of parameters within redundant layers to mitigate the performance drop caused by layer removal, these retained components are extremely low-rank and incur negligible inference overhead while preserving task performance.
+Additionally, we measure the compression time of GRASP and other structured pruning baselines when compressing LLaMA2-7B on an NVIDIA A100 GPU under a 25% compression ratio.
+As shown in Table 4, GRASP is able to compress the model efficiently while maintaining strong downstream performance, demonstrating its practicality for real-world deployment.
+Model Pruning Time (h) Compensation Time (h)
+LaCO 0.05 1.2
+SliceGPT 0.6 0.76
+LLM-Streamline 0.03 0.7
+GRASP 0.16 ×
+GRASP* 0.16 0.66
+Table 4: Compression time of GRASP and structured pruning baselines on LLaMA2-7B under a 25% compression ratio on a single A100 GPU.
+“×” indicates that GRASP does not require post-training compensation.
 Figure 5: Throughput of LLaMA2-7B and GRASP compressed model under 25% compression ratio on a single A100 GPU.
@@
 3.4 Ablation Study
-Calibration Data
-We conducted ablation studies to investigate the impact of representation evaluation dataset and the number of data used for singular value gradient attribution.
-Table 6 summarizes GRASP’s performance when compressing LLaMA3.1-8B-Instruct using different representation evaluation datasets and varying numbers of data from WikiText-2.
-The results demonstrate that GRASP is robust to the choice of both the representation evaluation dataset and the data number.
-Notably, GRASP achieves competitive performance with just 64 samples, highlighting its ability to perform effective compression with minimal data.
-Detailed results can be found in Appendix A.9.
+Calibration Data.
+We conduct ablation studies to assess the sensitivity of GRASP to the choice and size of the calibration dataset used for gradient-based singular value attribution.
+Table 5 and Figure 4 present the results on LLaMA3.1-8B-Instruct when varying the calibration dataset and the number of samples drawn from WikiText-2.
+The results indicate that GRASP remains robust across different calibration dataset choices and performs reliably even with limited calibration data.
+Remarkably, the method achieves strong performance with as few as 64 samples, demonstrating its effectiveness in low-data regimes.
 Ablation Type
@@
 512 37.86 63.97 47.12
-Table 6: Comparison of GRASP using different types and amounts of data for compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
+Table 5: Comparison of GRASP using different types and amounts of data for compressing LLaMA3.1-8B-Instruct at a 20% compression ratio.
+Results are reported on the WikiText-2, PTB datasets (measured by $\downarrow$) and the average accuracy across seven commonsense reasoning benchmarks: OpenBookQA, WinoGrande, HellaSwag, ARC-easy, ARC-challenge, PIQA, and MathQA.
 One-shot Pruning vs Iterative Pruning
-As detailed in Section 3.4, GRASP employs a layer-wise backward compression strategy to compress redundant layers.
+As detailed in Section 2.3, GRASP processes the redundant layers in a sequential manner, starting from the final redundant layer and proceeding backward through the network.
 This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers.
@@
 4 Discussion
-GRASP preserves representation stability.
-Layer removal disrupts the internal spectral structure of LLMs, often introducing inconsistencies in the learned representations.
-These disruptions lead to degradation in perplexity and loss of linguistic coherence, as the model struggles to compensate for missing transformations.
-GRASP stabilizes the compression process by selectively retaining representation-critical singular values, ensuring that the model’s residual stream remains structurally consistent.
-Unlike direct layer removal, which forces the model to undergo self-repair (Rushing and Nanda, 2024), GRASP proactively mitigates this instability by preserving the most influential spectral components, thereby reducing the burden on later layers to reconstruct lost information.
-GRASP accelerates adaptation via low-rank learning.
-Fine-tuning pruned models using parameter-efficient methods such as LoRA often introduces intruder dimensions into the pre-trained singular spectrum (Shuttleworth et al., 2024), leading to spectral dilution that weakens the contribution of pre-trained features.
-Similarly, replacing layers with dense modules, as in LLM-Streamline (Chen et al., 2024), requires training high-dimensional structures from scratch, which increases the risk of overfitting when data is limited.
-In contrast, GRASP identifies and fine-tunes the most essential low-rank subspaces within redundant layers, preserving the original spectral alignment while adapting efficiently to new tasks.
-Prior research has established the effectiveness of low-rank learning in deep networks (Oymak et al., 2019; Kamalakara et al., 2022), and GRASP leverages this principle to optimize task-specific performance while maintaining pre-trained feature integrity.
+The effectiveness of GRASP stems from two key design principles that distinguish it from conventional structured pruning methods.
+Leveraging Low-Rank Redundancy.
+GRASP builds on the observation that the redundant layers in LLMs exhibit inherent low-rank characteristics.
+Rather than removing these layers entirely, which can lead to information loss and degraded performance, GRASP retains a small subset of critical singular directions identified through a gradient-based attribution mechanism.
+This approach allows the model to retain essential functional capacity while discarding non-informative parameters, leading to efficient yet accurate compression.
+Preserving Informative Subspaces for Fast Recovery.
+By maintaining only the most influential parameters within redundant layers, GRASP avoids the need for costly retraining while facilitating faster convergence when post-training compensation is applied.
+This contrasts with approaches like LLM-Streamline (Chen et al., 2024) that insert randomly initialized or dense lightweight modules, which require more data and training time to approximate the original function.
+GRASP’s retention of critical low-rank components allows it to preserve the spectral alignment of the original model (Oymak et al., 2019, Kamalakara et al., 2022, Shuttleworth et al., 2024), resulting in more stable optimization dynamics and enhanced sample efficiency during post-training compensation.
+5 Related Work
+Structured pruning aims to reduce the size and computational cost of large language models (LLMs) by removing entire components such as layers, neurons, or dimensions, while maintaining model performance.
+Among this, layer pruning is a kind of structured pruning technique that eliminates redundant layers within Large Language Models (LLMs).
+Methods such as ShortGPT (Men et al., 2024) introduce a metric called Block Influence to assess the significance of individual layers, enabling efficient one-shot removal of less important layers.
+SLEB (Song et al., 2024) improves this by employing an iterative pruning strategy, evaluating the importance of each layer based on the current state of the layer-removed LLMs.
+LaCo (Yang et al., 2024), on the other hand, proposes a gradual compression approach, progressively merging redundant layers from deeper to shallower parts of the network.
+Although effective in reducing model size, layer pruning disrupts representation coherence, leading to performance degradation and increased perplexity, as analyzed by Liang et al.
+(2024).
+To mitigate this, post-training compensation methods have been proposed.
+Kim et al., 2024 introduced Shortened LLaMA, which employs LoRA (Hu et al., 2022) to restore pruned models’ capabilities.
+However, LoRA modifies the singular value spectrum, potentially weakening pre-trained features (Shuttleworth et al., 2024).
+LLM-Streamline (Chen et al., 2024) addresses this by training a lightweight module, such as an FFN or transformer layer, to approximate the removed layers.
+While effective, these methods impose high computational and data costs, limiting feasibility in resource-constrained settings.
+Another line of structured pruning research focuses on low-rank approximation, where Singular Value Decomposition (SVD) is widely used in model compression to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss.
+Recent methods have enhanced SVD to reduce LLM compression error.
+FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD.
+ASVD (Yuan et al., 2023) uses activation patterns from a calibration dataset to scale weight matrices, reducing compression-induced activation errors.
+SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss.
+Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices.
+In contrast to these approaches, GRASP integrates gradient-based attribution into the low-rank decomposition process, enabling both efficient compression and critical parameter retention.
 6 Conclusion
-We present GRASP, a novel framework for compressing large language models by selectively retaining representation-critical singular value directions.
-By mitigating the disruptions caused by layer removal, GRASP preserves essential transformations necessary for maintaining model stability and generalization.
-Experimental results demonstrate that GRASP achieves efficient and robust compression, improving perplexity and downstream task performance compared to existing methods while maintaining minimal computational overhead.
+In this work, we proposed GRASP, a novel compression framework that replaces redundant layers in large language models with a small set of adaptively selected singular parameters.
+By leveraging the low-rank structure of redundant layers and incorporating gradient-based attribution, GRASP identifies critical components that preserve model functionality with minimal parameter overhead.
+It operates in a training-free manner and enables efficient post-training recovery with limited data.
+Extensive experiments across diverse LLM architectures and benchmarks demonstrate that GRASP consistently outperforms existing structured pruning methods in both accuracy and efficiency.
 7 Limitations
-While GRASP demonstrates strong performance across various tasks and architectures, we acknowledge certain limitations that highlight opportunities for future research:
-Dependence on Calibration Data GRASP relies on representation evaluation dataset for singular value gradient attribution.
-In our experiments, using as few as 64 samples from the WikiText dataset achieved strong performance across multiple tasks.
-However, while this approach works well in these settings, we have not yet explored whether using domain-specific dataset could yield better results in specialized domains.
-Further testing with such data may provide valuable insights into its potential advantages.
-Trade-Off Between Compression and Retained Parameters To ensure robustness and task consistency, GRASP retains a small fraction of low-rank parameters in redundant layers, mitigating disruptions caused by pure layer removal.
-This design choice introduces a slight trade-off in terms of parameter retention, which could limit its suitability for applications requiring extreme compression ratios.
-Further reduction techniques may complement GRASP in such scenarios.
-Scalability to Extremely Large Models While GRASP performs effectively on models like LLaMA2 and LLaMA3, its application to ultra-large-scale models (e.g., hundreds of billions of parameters) could pose additional computational challenges.
-Future work could focus on optimizing the algorithm for such architectures, enabling its application to next-generation LLMs.
+While GRASP achieves competitive performance and compression efficiency, it also has several limitations that merit further exploration.
+First, our method relies on the assumption that layer redundancy can be identified via output similarity (e.g., cosine similarity between hidden states).
+While effective in practice, this heuristic may overlook more nuanced forms of redundancy that arise from distributed or task-specific behaviors within deeper model layers.
+Second, GRASP depends on access to gradient information and a small calibration dataset to compute attribution scores.
+Although the data requirement is minimal and the method is training-free in its core form, this may limit applicability in strictly black-box or privacy-sensitive settings where gradients or internal representations are inaccessible.
+We also note that our experiments are conducted on models up to 13B parameters and primarily in English-language tasks.
+Extending GRASP to multilingual or much larger-scale models is a promising direction for future work, especially as the scale and diversity of LLMs continue to grow.
 Ethical Considerations
@@
 $\frac{\partial\sigma_{i}}{\partial W}=u_{i}v_{i}^{T}$
-For a representation evaluation dataset $D$, the gradient of a singular value $\sigma_{i}$ with respect to the task loss can be interpreted as the projection of the weight gradient matrix $G$ onto the corresponding singular direction, given by:
+For a calibration dataset $D$, the gradient of a singular value $\sigma_{i}$ with respect to the task loss can be interpreted as the projection of the weight gradient matrix $G$ onto the corresponding singular direction, given by:
 $\frac{\partial L}{\partial\sigma_{i}}=u_{i}^{T}\frac{\partial L}{\partial W}v_{i}$
@@
 $\frac{\partial L}{\partial\Sigma}=I_{k}\odot U^{T}\left(\frac{\partial L}{\partial W}\right)V$
-A.2 Layer-wise Low-Rank Decomposition
-This section provides the implementation of Layer-wise Low-Rank Decomposition, as detailed in Algorithm 1.
-Algorithm 1 Gradient-Guided SVD for LLM Compression
-1:Model $M$ with $N$ layers, Representation Evaluation Dataset $D$, Retain Ratio $r$
-2:$\text{W}\in\{W_{q},W_{k},W_{v},W_{o},W_{up},W_{down}\}$ for weight matrices in self-attention and MLP blocks
-3:Compressed Model $\tilde{M}$
-4:Step 1: Redundant Layer Selection
-5:for $i=1$ to $N$ do
-6:     $\text{cos}_{i}\leftarrow\frac{H_{i}^{T}H_{i+1}}{\|H_{i}\|_{2}\|H_{i+1}\|_{2}}$
-7:end for
-8:Select $L$ layers with highest cosine similarity.
-9:Step 2: Gradient-guided SVD Compression:
-10:for each layer $l\in L$, starting from the last layer do
-11:     $\text{W}\in\{W_{q},W_{k},W_{v},W_{o},W_{up},W_{down}\}$
-12:     $\text{W}\leftarrow\mathbf{U}\Sigma\mathbf{V^{T}}$
-13:     $\text{G}\leftarrow\frac{\partial\mathcal{L}_{\text{}}}{\partial W}$   # compute gradient matrices using $D$
-14:     $\frac{\partial\mathcal{L_{\text{}}}}{\partial\sigma_{i}}\leftarrow\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$   # Compute gradients for singular values
-15:     $\tilde{W}\leftarrow\tilde{U}\tilde{\Sigma}\tilde{V^{T}}$   # Truncate singular values based on Eq.6, retain only $r\%$ parameters per matrix
-16:end for
-17:return $\tilde{M}$
+A.2 Experimental Setup and Hyperparameters Configuration
+To ensure a fair comparison, all experimental setups are consistent across all methods. In the following, we describe the experimental setup and hyperparameters configuration in detail.
+Hyperparameters Configurations
+For post-training compensation, all models compressed by GRASP are trained on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32.
+We use AdamW (Loshchilov and Hutter, 2019) as our optimizer and set the learning rate to $3\times 10^{-4}$.
+All our experiments are conducted on a single A100 GPU with mixed precision enabled.
+Table 6 provides the detailed configurations of post-training compensation.
+HyperParameters Setting
+Dataset Alpaca
+Huggingface Dataset Path yahma/alpaca-cleaned
+Batch Size 32
+Micro Batch Size 4
+Epochs 1
+Learning Rate 3.00E-04
+Max Length 256
+Train on Inputs TRUE
+Add EOS Token FALSE
+LoRA-Rank 256
+LoRA-Alpha 16
+LoRA-Dropout 0.05
+LoRA-Target-Modules
+q_proj, k_proj, v_proj, o_proj,
+up_proj, down_proj, gate_proj
+Prompt-Template Alpaca Template
+Table 6: Experimental setup and hyperparameters configurations.
 Method Metric Calibration Data Need Post-Training Training Data Training Dataset Size Training Module
@@
 Shortened LLaMA consists of two training stages: initial continual pre-training on the SlimPajama dataset, followed by LoRA fine-tuning on the Alpaca dataset.
-A.3 Experimental Setup and Hyperparameters Configuration
-To ensure a fair comparison, all experimental setups are consistent across all methods. In the following, we describe the experimental setup and hyperparameters configuration in detail.
-Metric
-For the WikiText-2 and PTB datasets, we use perplexity ($\downarrow$) to assess model’s generative capacity.
-For the seven common sense reasoning datasets, performance is measured by accuracy ($\uparrow$), and for the GSM8K dataset, performance is evaluated using Exact Match Accuracy ($\uparrow$).
-All common sense reasoning tasks are tested in zero-shot setting using the LM-Evaluation-Harness framework (Gao et al., 2024), while GSM8K is tested in the five-shot setting.
-Hyperparameters Configurations
-For post-training compensation, all models compressed by different methods are trained on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32.
-We use AdamW (Loshchilov and Hutter, 2019) as our optimizer and set the learning rate to $3\times 10^{-4}$.
-All our experiments are conducted on a single A100 GPU with mixed precision enabled.
-Table 8 provides the detailed configurations of post-training compensation.
-HyperParameters Setting
-Dataset Alpaca
-Huggingface Dataset Path yahma/alpaca-cleaned
-Batch Size 32
-Micro Batch Size 4
-Epochs 1
-Learning Rate 3.00E-04
-Max Length 256
-Train on Inputs TRUE
-Add EOS Token FALSE
-LoRA-Rank 256
-LoRA-Alpha 16
-LoRA-Dropout 0.05
-LoRA-Target-Modules
-q_proj, k_proj, v_proj, o_proj,
-up_proj, down_proj, gate_proj
-Prompt-Template Alpaca Template
-Table 8: Experimental setup and hyperparameters configurations.
-A.4 Comparison of Concurrent Pruning-based Methods
-We provide a detailed comparison of concurrent pruning-based LLM compression methods, and the results are summarized in Table 7.
-A.5 Detailed Results on Common Sense Reasoning Benchmarks
-In this section, we provide the detailed results for each benchmark.
-Table 9 summarizes the results of pruning-based methods without post-training compensation on the common sense reasoning benchmark, while Table 10 presents the results of methods with post-training compensation.
-Figure 4: Validation loss curves of different methods on the Alpaca dataset (1 epoch)
-Model Method Ratio Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
+A.3 Comparison of Concurrent Structured Pruning Methods
+We provide a detailed comparison of concurrent structured pruning LLM compression methods, and the results are summarized in Table 7.
+A.4 Detailed Results on Commonsense Reasoning Benchmarks
+In this section, we provide detailed results for GRASP and baseline methods on the commonsense reasoning benchmarks using LLaMA2-7B, LLaMA2-13B, and Mistral-7B.
+These results extend the main experiments presented in Table 1, offering a comprehensive view of GRASP’s performance across different model architectures.
+Table 8 reports accuracy without post-training compensation. The results demonstrate that GRASP consistently maintains strong accuracy and demonstrates robustness across model scales and families.
+A.5 More Results on Other Models
+To further validate the generalizability of GRASP across diverse LLM architectures, we present additional results under a 20% compression ratio on three representative models beyond LLaMA-7B, LLaMA2-13B, LLaMA3.1-8B-Instruct, and Mistral-7B.
+Table 9 summarizes the performance comparison between GRASP and structured low-rank pruning baselines, including FWSVD, ASVD, and SVD-LLM, across eight evaluation benchmarks.
+Consistent with our findings in the main text, GRASP achieves superior or comparable accuracy across most benchmarks.
+While SVD-LLM slightly outperforms GRASP on LLaMA2-13B, our method demonstrates stronger robustness across model families and benefits from more stable post-compensation recovery.
+These results highlight GRASP’s effectiveness as a generalizable and architecture-agnostic compression strategy.
+Model Method Ratio Openb. ARC_e WinoG. HeSW ARC_c PIQA MathQA Average
 Mistral-7B-v0.1 Dense 0.0% 0.33 0.81 0.74 0.61 0.50 0.81 0.36 0.59
@@
 Bold values indicate the best performance.
-Model Method Ratio Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
-Mistral-7B Dense 0.0% 0.33 0.81 0.74 0.61 0.50 0.81 0.36 0.59
-LoRA 25.1% 0.25 0.69 0.71 0.51 0.40 0.72 0.26 0.51
-LS-FFN 25.2% 0.22 0.67 0.68 0.50 0.38 0.70 0.27 0.49
-LS-Layer 24.1% 0.22 0.69 0.70 0.45 0.39 0.72 0.27 0.49
-GRASP 24.3% 0.26 0.69 0.70 0.52 0.41 0.73 0.27 0.51
-LLaMA-2-7B Dense 0.0% 0.32 0.69 0.67 0.57 0.40 0.78 0.28 0.53
-LoRA 25.0% 0.26 0.58 0.63 0.48 0.33 0.71 0.24 0.46
-LS-FFN 25.7% 0.25 0.56 0.63 0.48 0.33 0.71 0.25 0.46
-LS-Layer 24.0% 0.25 0.60 0.62 0.50 0.35 0.73 0.24 0.47
-GRASP 24.3% 0.28 0.61 0.65 0.50 0.36 0.72 0.25 0.48
-LLaMA-2-13B Dense 0.0% 0.32 0.73 0.70 0.60 0.46 0.79 0.30 0.56
-LoRA 25.0% 0.27 0.63 0.67 0.53 0.38 0.74 0.25 0.50
-LS-FFN 25.4% 0.28 0.61 0.67 0.54 0.38 0.73 0.25 0.50
-LS-Layer 24.4% 0.29 0.62 0.67 0.54 0.40 0.74 0.26 0.50
-GRASP 24.2% 0.29 0.63 0.68 0.55 0.39 0.74 0.27 0.51
-LLaMA3.1-8B-Instruct Dense 0.0% 0.34 0.82 0.74 0.59 0.52 0.80 0.39 0.60
-LoRA 22.3% 0.28 0.70 0.69 0.51 0.41 0.73 0.30 0.52
-LS-FFN 22.0% 0.26 0.70 0.70 0.49 0.41 0.72 0.30 0.51
-LS-Layer 21.7% 0.27 0.74 0.69 0.51 0.43 0.73 0.32 0.53
-GRASP 21.0% 0.28 0.72 0.70 0.52 0.43 0.74 0.31 0.53
-Table 10: Performance comparison of GRASP and pruning-based with post-training baselines under 25% compression ratio.
-LS-FFN and LS-Layer denote two variants of LLM-Streamline.
-Results are reported on seven reasoning datasets (individual and average accuracy).
-Bold values indicate the best performance.
-A.6 Validation Loss Curves
-We plot the validation loss curves when training pruned Mistral-7B with 20% compression ratio with different compensation methods.
-As shown in Figure 4, we can observe that GRASP has better initialization, leading to more stable and faster convergence during training.
-A.7 Evaluation Results on LongBench
+Mistral-7B LLaMA2-7B LLaMA2-13B LLaMA3.1-8B-Instruct
+Method
+PPL $\downarrow$
+(WikiText-2)
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+Acc $\uparrow$
+Average
+PPL $\downarrow$
+(WikiText-2)
+Acc $\uparrow$
+Average
+Original 5.25 0.59 5.68 0.52 5.47 0.53 7.21 0.60
+FWSVD 6357 0.32 1727 0.32 2360 0.31 3256.7 0.29
+ASVD 19.28 0.4 11.14 0.44 9.70 0.46 2443.99 0.30
+SVD-LLM 10.21 0.41 7.94 0.44 8.50 0.47 - -
+Ours 18.42 0.45 14.79 0.44 16.12 0.44 37.86 0.47
+Ours* 11.62 0.51 10.19 0.47 9.59 0.48 14.13 0.53
+Table 9: Perplexity($\downarrow$) of GRASP and low-rank pruning baselines on the WikiText-2 datasets and the average accuracy($\uparrow$) on seven common sense reasoning datasets for four different LLMs under 20% compression ratio.
+"†" indicates that we refer to the results in the original paper.
+The best performance is marked in bold.
+A.6 Evaluation Results on LongBench
 In this section, we present the detailed results of LLaMA3.1-8B-Instruct and its compressed version by GRASP under 20% compression ratio on LongBench, which are presented in Table 11.
 The results illustrate that GRASP with post-training compensation still maintains superior performance on long-form reasoning and complex generative tasks.
-A.8 Compression Costs of GRASP
-To assess the compressed costs of GRASP, we measure its compression time and peak memory consumption when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio.
-These metrics, summarized in Table 11, indicate that GRASP achieves a good balance between compression time and memory consumption.
-Model
-Average
-Compression Time(h)
-Compression Memory
-Consumption(MiB)
-LaCO 0.05 45046
-ASVD 5.5 17632
-SliceGPT 0.6 18195
-GRASP 0.16 29050
-Table 11: Compression time and peak compression memory usage of GRASP and baselines for LLaMA2-7B under 25% compression ratio on a single A100 GPU.
-A.9 Robustness of GRASP towards different Representation Evaluation Dataset
-In this section, we provide details of the ablation studies conducted to investigate the impact of representation evaluation datasets and the amount of data used for singular value gradient attribution.
+A.7 Robustness of GRASP towards different calibration Dataset
+In this section, we provide details of the ablation studies conducted to investigate the impact of calibration datasets and the amount of data used for singular value gradient attribution.
 Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of GRASP when compressing LLaMA3.1-8B-Instruct under 20% compression ratio.
@@
 All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage.
-As shown in Figure 5, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the representation evaluation dataset and the number of data.
-Tables 14 and 15 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different representation evaluation datasets (WikiText-2, C4) and varying numbers of calibration data.
-A.10 Robustness of GRASP towards Different Pruning Strategy
+As shown in Figure 4, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the calibration dataset and the number of data.
+Tables 12 and 13 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
+A.8 Robustness of GRASP towards Different Pruning Strategy
 In this section, we present the results of ablation experiments conducted with different pruning strategies in GRASP, with the results summarized in Table 10
@@
 The datasets are grouped as follows: (1-1 to 1-4) denote GovReport, QMSum, MultiNews, and VCSUM; (2-1 to 2-4) denote TREC, TriviaQA, SAMSum, and LSHT; (3-1 to 3-3) denote PassageCount, PassageRetrieval-en, and PassageRetrieval-zh; (4-1 to 4-4) denote NarrativeQA, Qasper, MultiFieldQA-en, and MultiFieldQA-zh; (5-1 to 5-4) denote HotpotQA, 2WikiMultihopQA, MuSiQue, and DuReader; (6-1 to 6-2) denote LCC and RepoBench-P.
-Figure 5: Performance of GRASP on LLaMA3.1-8B-Instruct under 20% compression using (a) different representation evaluation datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2.
-GRASP demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
-Representation Evaluation Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
+Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HeSW ARC_c PIQA MathQA Average
 WikiText-2 37.86 63.97 21.6 59.85 70.48 44.21 37.12 68.66 27.94 47.12
 C4 40.54 71.42 24 57.91 67.72 42.11 36.69 67.25 27.5 46.17
-Table 14: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression using 512 samples from WikiText-2 and C4 as representation evaluation datasets.
-Representation Evaluation Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
+Table 12: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression using 512 samples from WikiText-2 and C4 as calibration datasets.
+Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HeSW ARC_c PIQA MathQA Average
 WikiText-2-64 46.5 86.51 22.6 59.97 69.3 44.36 37.29 68.5 27.37 47.06

### v4

@@
 Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods, achieving 90% of the original model’s performance under 20% compression ratio.
-The source code is available at GRASP.
+The source code is available at https://github.com/LyoAI/GRASP.
 […]
@@
 Key Design.
-To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value based on its contribution to model performance, rather than relying on magnitude alone.
+To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value by its contribution to model performance, rather than relying on magnitude alone.
 Formally, the singular value decomposition of a weight matrix $W\in\mathbb{R}^{m\times n}$ is given by $W=U\Sigma V^{\top}$.
-Equivalently, $W$ can be represented as the sum of its rank-one components:
+Equivalently, $W$ can be expressed as the sum of its rank-one components:
 $W=\sum_{k=1}^{l}u_{k}\sigma_{k}v_{k}^{\top},$ (2)
@@
 $I(\Phi_{k})=T(\sigma_{k})+\sum_{i=1}^{m}T(u_{k,i})+\sum_{i=1}^{n}T(v_{k,i})$ (3)
-where $T(\cdot)$ denotes the estimated change in loss when a certain parameter $\theta$ is zeroed out.
-This can be approximated by the second-order Taylor series expansion (LeCun et al., 1989), which is defined as:
+where $T(\cdot)$ denotes the estimated loss change when a certain parameter $\theta$ is zeroed out. This can be approximated by the second-order Taylor expansion (LeCun et al., 1989), defined as:
 $T(\theta)=\left|\theta^{\top}\nabla_{\theta}\mathcal{L}+\frac{1}{2}\theta^{\top}H\theta+\mathcal{O}(\|\theta\|^{3})\right|$ (4)
 $\mathcal{L}=-\sum_{t}\log P(y_{t}\mid x_{\leq t})$ (5)
-where $\nabla_{\theta}\mathcal{L}$ denotes the gradient of the standard language modeling objective function $\mathcal{L}$ with respect to the parameter $\theta$, and $H$ represents the corresponding Hessian matrix.
+where $\nabla_{\theta}\mathcal{L}$ denotes the gradient of the standard language modeling objective function $\mathcal{L}$ w.r.t.
+the parameter $\theta$, and $H$ represents the corresponding Hessian matrix.
 To reduce computational overhead, we omit the second-order term and approximate the importance using only the first-order derivative (Hua et al., 2025, Kim et al., 2024):
@@
 The first term in Eq.
-6 captures the gradient projection of the loss $\mathcal{L}$ with respect to the singular value $\sigma_{i}$.
-The detailed proof is provided in Appendix A.1.
-By aggregating the first-order expansions across all components within each singular group, our method effectively captures the contribution of singular values.
-A larger value indicates a greater importance of the singular group.
+6 captures the gradient projection of the loss $\mathcal{L}$ w.r.t.
+the singular value $\sigma_{i}$.
+The proof is given in AppendixA.1.
+By aggregating the first-order expansions across all components within each singular group, our method effectively captures the contribution of singular values, where larger values indicate greater importance.
 This gradient-oriented attribution moves beyond heuristic magnitude-based criteria, enabling performance-aligned importance evaluation.
-Under the hypothesis that redundant layers exhibit an inherent low-rank structure, we retain only the top-$k$ singular values most critical to model performance based on Eq.
+Under the hypothesis that redundant layers exhibit an inherent low-rank structure, we retain only the top-rr% singular values most critical to model performance based on Eq.
 6, and use them to replace the corresponding redundant layer.
+The relationship between the singular group retain ratio (rr%) and the overall target compression ratio is detailed in Appendix A.3.
 2.3 Detailed Implementation of GRASP
@@
 3:Step 1: Redundant Layer Selection:
-4:Compute cosine similarity $\text{cos}(H_{i},H_{i+1})$ for all layers
+4:Compute cosine similarity $\text{cos}(H_{i},H_{i+1})$ for all layers via Eq.1 using $D$
 5:Select top-$L$ layers with highest similarity as redundant
@@
 7:for each redundant layer $l$ (in reverse order) do
-[…]
+8:  for each $W\in\{$attention, MLP}\} do
 9:   SVD: $W=U\Sigma V^{\top}$
-10:         Compute importance $I(\Phi_{k})$ for each $\Phi_{k}=\{u_{k},\sigma_{k},v_{k}^{\top}\}$ via Eq. 6
+10:   Compute importance $I(\Phi_{k})$ for each $\Phi_{k}=\{u_{k},\sigma_{k},v_{k}^{\top}\}$ via Eq.6 using $D$
 11:   Keep top-$r\%$ singular groups, reconstruct $\tilde{W}$
@@
 GRASP 44.6 35.1 26.2 68.4 41.4 63.2 62.7 73.3 35.1 36.1 43.1 30.7 46.7 94.9%
-Table 2: Comparison between GRASP and structured pruning baselines with post-training compensation under a 20% compression ratio.
+Table 2: Comparison between GRASP and structured pruning baselines with post-training compensation under a 25% compression ratio.
 Results marked with * are reported from (Chen et al., 2024).
@@
 3.2 Comparison with Pruning-based LLM Compression Methods
-3.2.1 Experiment 1
+3.2.1 Comparison without Post-Training Compensation
 Models and Benchmarks.
@@
 Notably, GRASP demonstrates significantly improved robustness across models, effectively mitigating the variability in pruning sensitivity across diverse architectures.
-3.2.2 Experiment 2
+3.2.2 Comparison with Post-Training Compensation
 Models and Benchmarks.
@@
 In addition to achieving superior accuracy, GRASP also demonstrates more stable and faster convergence during post-training, which we attribute to its preservation of critical singular components.
-3.2.3 Experiment 3
+3.2.3 Comparison with Low-Rank Pruning Methods
 Models and Benchmarks.
@@
 Additionally, we measure the compression time of GRASP and other structured pruning baselines when compressing LLaMA2-7B on an NVIDIA A100 GPU under a 25% compression ratio.
-As shown in Table 4, GRASP is able to compress the model efficiently while maintaining strong downstream performance, demonstrating its practicality for real-world deployment.
+As reported in Table 4, Pruning Time refers to the time required to perform model compression up to the target sparsity, excluding any compensation.
+Compensation Time refers to the time spent on post-compression procedures—such as fine-tuning or parameter updates—aimed at recovering model performance.
+The results show that GRASP is able to compress the model efficiently while maintaining strong performance, demonstrating its practicality for real-world deployment.
 Model Pruning Time (h) Compensation Time (h)
@@
 512 37.86 63.97 47.12
-Table 5: Comparison of GRASP using different types and amounts of data for compressing LLaMA3.1-8B-Instruct at a 20% compression ratio.
-Results are reported on the WikiText-2, PTB datasets (measured by $\downarrow$) and the average accuracy across seven commonsense reasoning benchmarks: OpenBookQA, WinoGrande, HellaSwag, ARC-easy, ARC-challenge, PIQA, and MathQA.
+[…]
+Retain Ratio rr%.
+We further explore the impact of the retain ratio $r$ on GRASP’s performance.
+As shown in Table 6 and Figure 6, performance improves rapidly when increasing $r$ from 0% to 10%, after which the gains largely saturate.
+This indicates that a retain ratio of 10% serves as a practical inflection point, providing a strong trade-off between compression and accuracy.
+In practice, we therefore adopt a fixed retain ratio of 10%, which consistently yields stable performance across benchmarks while maintaining high compression efficiency.
+These results further confirm that the selected redundant layers contain substantial redundancy, as even a small fraction of retained singular components suffices to recover most of the original performance.
+Retain Ratio
+Overall
+Compression Ratio
+Openb. ARC_e ARC_c WinoGrande PIQA MathQA Average
+0% 24.0% 20.2 45.1 30.9 61.3 67.0 22.2 41.1
+5% 22.8% 22.4 54.9 31.8 66.0 70.0 23.8 44.8
+10% 21.6% 23.2 56.1 32.7 66.2 70.7 23.9 45.5
+15% 20.4% 23.4 56.3 32.9 65.6 70.7 24.1 45.5
+20% 19.2% 23.4 56.5 33.4 65.6 71.5 24.3 45.8
+Table 6: Ablation results on LLaMA2-7B by replacing 8 redundant layers with sensitivity-aware singular parameters while varying the retain ratio.
+The table reports the resulting overall compression ratio and downstream task performance across multiple benchmarks.
+Figure 6: Effect of the retain ratio on average performance when compressing 8 redundant layers of LLaMA2-7B.
+The results show a sharp improvement up to a 10% retain ratio, after which performance gains plateau, indicating 10% as a practical inflection point.
 One-shot Pruning vs Iterative Pruning
@@
 This can be done using one-shot pruning, which identifies and decomposes all redundant layers in a single step, or iterative pruning, which processes layers one at a time to account for interactions between layers.
-We present the ablation results in Table 10 in Appendix A.8, which shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
+We present the ablation results in Table 7, which shows that both approaches achieve similar performance, with one-shot pruning being more efficient.
+Pruning Strategy WikiText-2 PTB
+Average
+Accuracy
+Compression
+Time(h)
+One-shot Pruning 37.86 63.97 47.12 0.16
+Iterative Pruning 38.39 72.18 47.13 0.22
+Table 7: Comparison of one-shot pruning and iterative pruning for LLaMA3.1-8B-Instruct under 20% compression ratio.
 4 Discussion
@@
 While effective, these methods impose high computational and data costs, limiting feasibility in resource-constrained settings.
-Another line of structured pruning research focuses on low-rank approximation, where Singular Value Decomposition (SVD) is widely used in model compression to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss.
+Another line of structured pruning research focuses on low-rank approximation, where Singular Value Decomposition (SVD) is widely used to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss.
 Recent methods have enhanced SVD to reduce LLM compression error.
@@
 SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss.
-Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices.
-In contrast to these approaches, GRASP integrates gradient-based attribution into the low-rank decomposition process, enabling both efficient compression and critical parameter retention.
+Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output activations with low-rank matrices.
+However, SVD-based methods require truncating at least 50% of singular values to reduce parameters of square matrices (which are common in LLMs like Llama), which often leads to significant information loss.
+In contrast to these approaches, GRASP integrates gradient-based attribution into the low-rank decomposition process and replaces redundant layers with only a small fraction of parameters (typically 10%), thereby enabling efficient compression while retaining critical information.
 6 Conclusion
@@
 Shortened LLaMA consists of two training stages: initial continual pre-training on the SlimPajama dataset, followed by LoRA fine-tuning on the Alpaca dataset.
-A.3 Comparison of Concurrent Structured Pruning Methods
+A.3 Relationship between Singular Group Retail Ratio and Target Overall Compression Ratio
+In GRASP, the retain ratio rr% denotes the proportion of singular components preserved within each weight matrix of the selected redundant layers.
+The target overall model compression ratio, which is jointly determined by (i) the number of layers selected in Step 1, and (ii) the retain ratio rr% applied in Step 2.
+In practice, we adopt a fixed retain ratio—typically 10%—as ablation studies (Section 3.4) show it achieves a favorable trade-off between compression and performance.
+Table 6 further illustrates how varying the retain ratio influences the overall compression ratio and downstream performance across several benchmarks when compressing 8 layers of LLaMA2-7B.
+These results highlight the high redundancy of the selected layers: even a small fraction of retained singular components is sufficient to recover most of the original performance.
+A.4 Comparison of Concurrent Structured Pruning Methods
 We provide a detailed comparison of concurrent structured pruning LLM compression methods, and the results are summarized in Table 9.
-A.4 Detailed Results on Commonsense Reasoning Benchmarks
+A.5 Detailed Results on Commonsense Reasoning Benchmarks
 In this section, we provide detailed results for GRASP and baseline methods on the commonsense reasoning benchmarks using LLaMA2-7B, LLaMA2-13B, and Mistral-7B.
@@
 Table 10 reports accuracy without post-training compensation. The results demonstrate that GRASP consistently maintains strong accuracy and demonstrates robustness across model scales and families.
-A.5 More Results on Other Models
+A.6 More Results on Other Models
 To further validate the generalizability of GRASP across diverse LLM architectures, we present additional results under a 20% compression ratio on three representative models beyond LLaMA-7B, LLaMA2-13B, LLaMA3.1-8B-Instruct, and Mistral-7B.
@@
 The best performance is marked in bold.
-A.6 Evaluation Results on LongBench
+A.7 Evaluation Results on LongBench
 In this section, we present the detailed results of LLaMA3.1-8B-Instruct and its compressed version by GRASP under 20% compression ratio on LongBench, which are presented in Table 12.
 The results illustrate that GRASP with post-training compensation still maintains superior performance on long-form reasoning and complex generative tasks.
-A.7 Robustness of GRASP towards different calibration Dataset
+A.8 Robustness of GRASP towards different calibration Dataset
 In this section, we provide details of the ablation studies conducted to investigate the impact of calibration datasets and the amount of data used for singular value gradient attribution.
@@
 As shown in Figure 4, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the calibration dataset and the number of data.
-Tables 12 and 13 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
-A.8 Robustness of GRASP towards Different Pruning Strategy
-In this section, we present the results of ablation experiments conducted with different pruning strategies in GRASP, with the results summarized in Table 10
-Pruning Strategy WikiText-2 PTB
-Average
-Accuracy
-Compression
-Time(h)
-One-shot Pruning 37.86 63.97 47.12 0.16
-Iterative Pruning 38.39 72.18 47.13 0.22
-Table 10: Comparison of one-shot pruning and iterative pruning for LLaMA3.1-8B-Instruct under 20% compression ratio.
+Tables 13 and 14 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
+A.9 Comparison with Unstructured Pruning Methods
+To provide a more comprehensive picture of model compression, we have included comparisons with two widely used unstructured pruning methods, SparseGPT (Frantar and Alistarh, 2023) and Wanda (Sun et al., 2023) in our experiments.
+While unstructured pruning introduces fine-grained sparsity in weight matrices, it typically requires dedicated sparse formats to realize storage or runtime benefits, which complicates direct comparison under the same compression ratio.
+To ensure a fair evaluation, we adopt the commonly used 2:4 sparsity pattern (i.e., 50% sparsity) for both methods, and compare them against GRASP at a 20% structured compression ratio.
+The results are summarized in Table 15, indicating that GRASP achieves competitive or superior performance against unstructured pruning methods across all benchmarks.
 Model Summarization Few-shot Learning Synthetic Task
@@
 Table 14: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression with varying calibration data sizes (64, 128, 256, 512) from WikiText-2.
+Method Openb. ARC_e ARC_c WinoG. HeSW PIQA MathQA Average
+Dense 32.12 72.69 39.87 67.33 57.16 78.46 26.02 53.38
+SparseGPT(2:4) 24.40 64.10 30.80 66.61 43.09 70.73 24.25 46.28
+wanda(2:4) 24.60 62.75 30.80 62.67 41.38 70.29 23.62 45.16
+GRASP 24.80 56.02 32.68 66.14 42.57 70.74 23.92 45.27
+GRASP* 28.20 68.73 37.63 67.88 50.98 72.47 24.19 50.01
+Table 15: Comparison between GRASP and unstructured pruning methods under the 2:4 sparsity pattern (50% sparsity).
+GRASP is applied at a 20% structured compression ratio.
+GRASP* denotes the variant with light performance compensation